### v1

# Beyond Conventional Transformers: The Medical X-ray Attention (MXA) Block for Improved Multi-Label Diagnosis Using Knowledge Distillation

Abstract
Medical imaging, particularly X-ray analysis, often involves detecting multiple conditions simultaneously within a single scan, making multi-label classification crucial for real-world clinical applications. In this paper, we present the Medical X-ray Attention (MXA) block, a novel attention mechanism tailored specifically to address the unique challenges of X-ray abnormality detection. The MXA block enhances traditional Multi-Head Self Attention (MHSA) by integrating a specialized module that efficiently captures both detailed local information and broader global context. To the best of our knowledge, this is the first work to propose a task-specific attention mechanism for diagnosing chest X-rays, as well as to attempt multi-label classification using an Efficient Vision Transformer (EfficientViT). By embedding the MXA block within the EfficientViT architecture and employing knowledge distillation, our proposed model significantly improves performance on the CheXpert dataset, a widely used benchmark for multi-label chest X-ray abnormality detection, with data augmentation methods applied. Our approach achieves an area under the curve (AUC) of 0.85, an absolute improvement of 0.19 compared to our baseline model’s AUC of 0.66, corresponding to a substantial approximate 233% relative improvement above random guessing (AUC = 0.5).  [^1] The code and results for this study are available at: https://github.com/Hadi-M-Ibrahim/Beyond-Conventional-Transformers.
Keywords Computer Vision  $\cdot$ Medical Imaging  $\cdot$ Multi-Label Classification  $\cdot$ Vision Transformers  $\cdot$ X-ray Diagnosis
1 Introduction
The adaptation of the attention-based transformer architecture [1] for computer vision, leading to the development of Vision Transformers (VITs) [2], demonstrates exceptional performance in tasks such as image classification [2], segmentation [3], and generation [4]. The ViT showcases superior model performance in contrast to traditional methodologies such as Recurrent and Convolutional Neural Networks [2]. However, these advances in accuracy came with the downside of an increased model size and computational cost.
Important breakthroughs have been made in medical analysis with the adoption of deep learning techniques, particularly CNNs [5] and transformers. CNN-based architectures like U-Net [6] have set benchmarks in segmentation tasks, enabling precise delineation of organs and anomalies in medical images. However, non-CNN architectures such as transformers and hybrid models (e.g., TransUNet [7]) have recently gained traction, leveraging the global attention capabilities of transformers to achieve superior performance in tasks like segmentation. For instance, Swin Transformers [8], with their hierarchical structure and shifted windowing mechanism, have been instrumental in improving the efficiency and scalability of vision models, making them suitable for tasks requiring high-resolution input. Additionally, models like CheXNet [9], trained on chest X-rays, have been pivotal in diagnosing diseases like pneumonia with performance comparable to radiologists.
Despite the advancements in transformer-based architectures for vision tasks, their practical deployment in clinical domains like medical imaging faces significant challenges. Transformers typically require extensive computational resources and large-scale datasets for training, both of which are often scarce in medical imaging [10, 7]. This limitation is particularly pronounced in X-ray analysis, where datasets are typically smaller, and many publicly available datasets suffer from issues such as inconsistent data collection procedures and biased performance assessment metrics, leading to poor generalization on unseen data [11, 12, 13]. These challenges underscore the need for models that are both computationally efficient and capable of handling the inherent data limitations in medical imaging.
Efforts to address these challenges have led to the development of efficient transformer architectures, such as Data-efficient Image Transformers (DeiTs) [14] and EfficientViTs (EfficientViTs) [15], which reduce computational overhead while maintaining competitive performance. For instance, DeiTs leverage knowledge distillation to improve efficiency, while EfficientViTs optimize the balance between Multi-Head Self-Attention (MHSA) and Feed-Forward Network (FFN) layers to achieve significant gains in throughput and accuracy. These advancements make them particularly suited for medical imaging applications, including X-ray analysis, where computational resources are limited and scalability is critical. The shift from CNNs to pure transformer architectures is driven by the state-of-the-art performance of Vision Transformers (ViTs) in vision tasks such as ImageNet classification [2], highlighting their immense potential contribution to the field of medical imaging.
Additionally, multi-label classification is a crucial yet underexplored area in X-ray diagnostics, as a single scan often reveals multiple conditions simultaneously [9, 16]. In this work, we propose a framework to address these gaps. Specifically, our contributions are as follows:
  1. 1.
    Multi-Label EfficientViTs: We adapt the EfficientViT (EfficientViT) model [15] to handle multi-label outputs. By replacing cross-entropy loss with BCEWithLogitsLoss and introducing CheXpert [16], our approach efficiently handles multiple pathologies within a single forward pass.
  2. 2.
    Medical X-ray Attention (MXA) Block: We introduce the MXA block, which integrates Dynamic Region of Interest (ROI) Pooling [17] and CBAM [18]-style attention in parallel with the standard Multi-Head Self-Attention layers. This aims to capture both localized abnormal findings (often subtle in chest X-rays) and global context without incurring large computational overhead.
  3. 3.
    Multi-label Knowledge Distillation: To further boost performance under data constraints, we employ a multi-label distillation scheme. A teacher model (e.g., DenseNet-121 [14] pre-trained on CheXpert [16]) provides soft or hard labels, thereby improving the student transformer’s recognition of difficult pathologies.
We validate our method on CheXpert, demonstrating that our multi-label EfficientViTs with the MXA block achieve improved accuracy and computational efficiency. Notably, the parallel integration of MHSA and MXA—augmented with knowledge distillation [19]—enables both robust global awareness and fine-grained local feature extraction, addressing common pitfalls in chest X-ray interpretation.
The rest of the paper is organized as follows: we review related works in Section 2, including advancements in multi-label image classification and medical imaging techniques. Section 3 details our approach to adapting EfficientViTs for multi-label classification. Section 4 introduces the Medical X-ray and Section 5 outlines our approach to Knowledge Distillation. Section 6 explains our experimental setup, including the dataset, evaluation metrics, and implementation details. Section 7 presents the results and the implications of our findings, limitations, and future directions for research in transformer-based medical imaging models.
2 Related Work
2.1 Multi-label Image Classification
Image classification has been a key application of computer vision, with substantial advancements driven by deep learning. Convolutional Neural Networks (CNNs) [5] have historically been the backbone of image classification, achieving state-of-the-art performance on datasets like ImageNet [20]. Architectures such as AlexNet [21], ResNet [22], and EfficientNet [10] have progressively improved model performance by introducing innovations in depth, width, and computational efficiency. More recently, Vision Transformers (ViTs) [2] have demonstrated the ability to outperform CNNs by leveraging self-attention mechanisms, particularly in large-scale datasets, thereby highlighting the potential of transformer-based architectures in image classification.
While significant progress has been made in single-label image classification, multi-label image classification presents additional challenges. Unlike single-label tasks, where each image is associated with a single category, multi-label classification requires models to predict multiple, potentially overlapping labels for a single image. This is particularly important in real-world applications such as clinical settings of medical imaging, where a single scan can reveal multiple conditions. Standard deep learning architectures, such as ResNet and DenseNet, have been adapted to multi-label classification by employing sigmoid non-linearity functions and binary cross-entropy loss to handle multiple labels independently [23, 24].
In the medical imaging domain, multi-label classification has been explored for tasks such as diagnosing chest X-rays, where an image may indicate multiple abnormalities. CheXNet [9] was among the first models to utilize a deep learning-based approach for chest X-ray classification, achieving radiologist-level performance on pneumonia detection while identifying other conditions. Later works, such as those leveraging the CheXpert dataset [16], highlighted the challenges of handling label uncertainty and imbalance in multi-label datasets. Transformers have recently been explored in this context due to their ability to model global dependencies, making them well-suited for identifying co-occurring conditions. For example, TransMed [25] applies transformer-based models for multi-label classification in the medical domain, demonstrating their potential in improving diagnostic accuracy.
2.2 CheXNet: Pneumonia Detection in Chest X-rays
CheXNet [9] is a pioneering work in the application of deep learning to chest X-ray analysis. This model utilizes a 121-layer DenseNet [26] architecture to perform multi-label classification, diagnosing up to 14 thoracic diseases. The dataset used for training was the NIH Chest X-ray dataset [27], which contains more than 100,000 images labeled with multiple conditions. The primary innovation of CheXNet is its ability to achieve radiologist-level performance on pneumonia detection, demonstrating the efficacy of CNNs [5] for X-ray diagnostics.
Despite its success, CheXNet is constrained by its reliance on large-labeled datasets, and its performance can degrade when faced with noisy or unbalanced data. Furthermore, while CheXNet addresses multi-label classification, it does not explicitly model dependencies between labels, which limits its ability to handle co-occurring conditions effectively.
Building on these limitations, we aim to leverage the power of transformer-based architectures to further improve performance in X-ray diagnostics. By combining methodologies such as knowledge distillation and cascading group attention, we aim to achieve higher accuracy in multi-label classification while introducing methods to reduce computational overhead.
2.3 CheXpert: A Benchmark for Chest X-ray Diagnosis
CheXpert [16] is a large-scale dataset designed to benchmark the performance of machine learning models in chest X-ray diagnosis. It consists of over 224,000 labeled chest radiographs from 65,000 patients, annotated with uncertainty labels for 14 common thoracic conditions. CheXpert addresses some of the challenges faced by prior datasets, such as label imbalance and inconsistency, by employing a rule-based labeling system to extract observations from radiology reports. This enables the creation of a standardized dataset that serves as a robust benchmark for evaluating machine learning models.
The CheXpert dataset has driven innovation in chext X-ray analysis. Current top-performing approaches often integrate advanced architectures, such as DenseNets and transformers, with tailored loss functions and training strategies. For example, in "Large-scale Robust Deep AUC Maximization" [28], a novel surrogate loss function is introduced designed to maximize the Area Under the Curve (AUC) for imbalanced datasets. This method, tested on CheXpert, achieves state-of-the-art results by directly optimizing AUC, a metric critical in medical imaging tasks.
Despite these advances, the computational cost associated with these models remains a challenge, particularly for real-time or resource-constrained applications. Our work builds upon these advancements by introducing EfficientViTs coupled with knowledge distillation, aiming to improve computational efficiency while maintaining state-of-the-art performance on benchmarks such as CheXpert.
2.4 DenseNet-121
The Dense Convolutional Network (DenseNet) architecture was proposed by Huang et al. [26] to improve information flow in convolutional neural networks by introducing dense connectivity patterns. Unlike traditional CNNs, where each layer only connects to the following layer, DenseNets connect each layer to every other layer in a feed-forward manner. This design drastically enhances gradient propagation and feature reuse throughout the network, enabling models to be both deeper and more parameter-efficient.
DenseNet-121 [26] specifically has 121 layers organized in dense blocks separated by transition layers, striking a good balance between accuracy and computational cost. In medical imaging, DenseNet-121 has been especially popular due to its ability to learn rich feature representations from relatively limited datasets. For instance, Rajpurkar et al. [9] employed a DenseNet-121 backbone in CheXNet—one of the pioneering works that applied deep learning to chest X-rays—achieving near-radiologist-level performance in detecting pneumonia. Subsequent adaptations of DenseNet-121, often fine-tuned on large chest radiograph datasets (e.g., NIH ChestX-ray14 [27] or CheXpert [16]), have shown state-of-the-art results for a variety of thoracic abnormalities.
3 Multi-labeling with Efficient Vision Transformers
While the original EfficientViTs [15] implementation was geared towards single-label tasks (e.g. ImageNet classification), we extend its functionality to multi-label classification, an essential capability in clinical X-ray diagnosis, where multiple abnormalities may coexist. This section details our modifications to EfficientViTs, including the transition from cross-entropy (CE) to multi-label loss, changes to the dataset loader, and the integration of a multi-label knowledge distillation configuration.
3.1 Adjusting the Loss Function and Output Layer
In single-label classification, the training loop typically relies on a CE loss [29] with softmax outputs. However, multi-label classification requires each sample to predict a set of labels independently. Therefore, we replace the single-label objective with Binary Cross Entropy with Logits (BCEWithLogitsLoss), which effectively combines a sigmoid activation and binary cross-entropy in one step.
Let $o=(o_{1},...,o_{C})$ be the raw logits output by the network for $C$ classes, and let $y=(y_{1},...,y_{C})$ be the multi-label ground truth, where each $y_{C}\in[0,1]$. BCEWithLogitsLoss is defined as:
$\mathscr{L}_{\text{BCEWithLogits}}(o,y)=-\sum_{c=1}^{C}[y_{c}\text{log}(\sigma(o_{c}))+(1-y_{c})\text{log}(1-\sigma(o_{c}))]$
where $\sigma$ is the sigmoid function. In our case, we set $C=14$ to match the CheXpert dataset [16]. After the final pooling layer, we thus produce logits of shape $(B,14)$ and if multi-labeling is active, we apply $\sigma$ to map logits to probabilities $[0,1]$ for each pathology.
3.2 Dataset and Data Loading for Multi-label Targets
In contrast to single-label datasets (e.g., ImageNet), our CheXpert dataset [16] provides annotations for up to 14 possible pathologies per image. We implement a custom CheXpertDataset class that reads the CheXpert CSV and constructs a multi-label vector for each X-ray. Specifically, for each row we parse 14 binary indicators (e.g., Atelectasis, Cardiomegaly, Pneumonia) and store them in a tensor of shape (14,). We follow the “U-1” protocol, treating uncertain (U) labels as positive (i.e., 1) rather than negative (0). This choice is motivated by findings in [16] and confirmed by Table 2 in [30], which show improved performance when uncertain labels are approximated as positive. Formally, for each pathology label:
$y_{c}=\begin{cases}1,&\text{if the label is 1 or uncertain},\\ 0,&\text{otherwise}.\end{cases}$
Though uncertain annotations can be mapped to 0, 1, or even ignored entirely, recent work suggests that “U-1” often yields better diagnostic performance than “U-0,” especially for pathologies where missing or subtle features might still be clinically relevant [30, 11]. Our approach is straightforward yet effective, and can be adapted depending on domain-specific requirements.
This design integrates well with the multi-label BCEWithLogitsLoss, allowing the network to learn from both definitively positive and uncertain/ambiguous labels in parallel. As discussed further in Section 3.3, we feed these labeled samples into the EfficientViT training loop, ensuring that each of the 14 possible pathologies can be predicted independently per image.
3.3 Training and Evaluation
Our training loop (adapted from EfficientViTs) still leverages the original DistillationLoss wrapper but replaces the base cross-entropy criterion with BCEWithLogitsLoss. Let $B$ be the batch size and let $O\in\mathbb{R}^{B\times 14}$ be the output logits. The multi-label ground truth is $Y\in[0,1]^{B\times 14}$. During each forward pass,
[…]
and the weights of EfficientViTs are updated via standard backpropagation.
During validation, we compute the same loss. Additionally, to track a simple measure of correctness for each label [31], we compute an element-wise comparison of $\sigma(O)$ with $Y$, thresholding at $0.5$. The resulting “accuracy” reflects the fraction of individual pathology labels predicted correctly across the batch.
3.4 Knowledge Distillation with a Multi-label Teacher
To further improve performance, especially for challenging or subtle pathologies, we apply knowledge distillation (KD) in a multi-label context. Our DistillationLoss class uses a teacher model (a pre-trained DenseNet-121 on CheXpert with 14 outputs) and compares the student’s predictions to the teacher’s logits transformed by a sigmoid. Formally, let $O^{s}$ by the student’s output and $O^{t}$ be the teacher’s output. TEH teacher probabilities are
$p^{t}=\sigma(O^{t})$
and we compute a distillation loss that is either:
  1. 1.
    Soft Distillation: Using the teacher probabilities directly in a BCE-like loss against the student logits:
    $\mathscr{L}^{\text{soft}}_{\text{distill}}=\text{BCEWithLogitsLoss}(O^{s},p^{t})$
  2. 2.
    Hard Distillation: Binarize the teacher probabilities at $0.5$ to form pseudo-labels $h^{t}=\mathbb{I}\bigl{[}p_{t}>0.5\bigr{]}$, then compute a BCE loss between $O^{s}$ and $h^{t}$.
In both cases, the total loss is a weighted combination of the base BCEWithLogitsLoss and $\mathscr{L}_{\text{distill}}$. Formally, if $\alpha$ is the distillation weight,
$\mathscr{L}_{\text{total}}=(1-\alpha)\mathscr{L}_{\text{BCEWithLogits}}(O^{s},Y)+\alpha\mathscr{L}_{\text{distill}}(O^{s},O^{t})$
Through this multi-label KD strategy, the student model more effectively learns to approximate the teacher’s predictions, leading to improved accuracy.
4 Medical X-ray Attention (MXA) Block
We propose a new block designed to enhance the efficiency and accuracy of transformer-based architectures in multi-label clinical X-ray abnormality detection and diagnosis. The MXA block is composed of Dynamic Region of Interest (ROI) [17] Pooling and Convolutional Block Attention Module (CBAM)-Style Attention [18], each of which focuses on improving localized model performance. We integrate this block in parallel with the Multi-Head Self-Attention (MHSA) layers within EfficientViTs [15] to optimize both computational efficiency and accuracy.
Figure 1: Integration of the Medical X-ray Attention (MXA) Block Architecture.
4.1 Dynamic Region-of-Interest (ROI) Pooling
Multi-label X-ray classification often includes multiple abnormal regions and purely global feature pooling might dilute small or subtle findings. To identify the most relevant regions for targeted analysis, consequently reducing computational overhead, we propose a Dynamic ROI pooling approach. By dynamically pooling region-of-interest (ROI) features, the model subsectionasizes critical areas while ignoring extraneous regions.
Given an input feature map $F\in\mathbb{R}^{BxCxHxW}$, where $B$ is the batch size, $C$ is the number of channels, $H,W$ are the height and width of the feature map, our ROI pooling mechanism selects critical regions using a region predictor, formally defined as:
$\text{ROI}_{i}=[x_{1},y_{1},x_{2},y_{2}],\ \text{where}\ i=1,2,\ldots,B$
Here, $[x_{1},y_{1}]$ and $[x_{2},y_{2}]$ are normalized coordinates of the top-left and bottom-right corners of the region of interest (ROI) for the $i$-th image. These coordinates are generated by a lightweight convolutional network conditioned on the feature map. The bounding-box region predictor is trained jointly with the classification objective in an end-to-end fashion: during the forward pass, each predicted ROI is used to crop and rescale the input feature map, and the resulting ROI features flow into the rest of the network for classification. The gradients from the multi-label classification loss then back-propagate through the ROI predictor as well, optimizing both the bounding boxes and the classification parameters simultaneously. Consequently, the network learns ROIs that best localize relevant abnormalities for improved multi-pathology detection.
After cropping the predicted ROI from the feature map, we resize it to match the original feature map size via bilinear interpolation:
$F^{(i)}_{\text{pooled}}=\text{Resize}(F_{\text{ROI}_{i}},(H,W))$
where $F^{(i)}_{\text{pooled}}$ is the pooled ROI feature map, and Resize()($\cdot$)( ⋅ ) denotes the bilinear interpolation function. By pooling only the relevant region to full resolution, we reduce the overhead that would otherwise come from processing the entire X-ray at high granularity [32]. This ensures compatibility with subsequent operations while focusing on the most critical regions. We use a kernel size of 3 in certain layers (e.g. local window attention or ROI-based convolutional components) as a practical middle ground for ROI between capturing sufficient local context and avoiding over-fitting in chext X-ray data. Prior work in medical image analysis (e.g., [6]) has similarly found that intermediate kernel sizes strike a good balance between detail preservation and robust feature extraction.
4.2 CBAM-Style Attention
To further refine the representation and enhance localized feature learning, we incorporate a Convolutional Block Attention Module (CBAM) [18] into the MXA block. CBAM sequentially applies channel attention and spatial attention to the pooled feature maps, enabling the network to focus on the most informative channels and spatial regions. This combined attention is especially helps the network prioritize more pathologically relevant features, crucial in medical images where anomalies can be small or faint. Spatial in particular helps improve the overall global context.
Channel attention determines which feature channels contribute the most to the classification task. This is achieved by globally aggregating the feature map along the spatial dimensions using global average pooling (GAP) and global max pooling (GMP), followed by a shared multi-layer perceptron (MLP) [33]. For a given pooled feature map $F_{pooled}\in\mathbb{R}^{C\times H\times W}$, channel attention is computed as:
$M_{C}=\sigma(W_{2}\delta(W_{1}\text{GAP}(F_{\text{pooled}}))+W_{2}\delta(W_{1}\text{GMP}(F_{\text{pooled}})))$
where $W_{1}\in\mathbb{R}^{C\times\frac{C}{r}}$ and $W_{2}\in\mathbb{R}^{\frac{C}{r}\times C}$ are learnable weight matrices, $\delta$ is the ReLU activation function [34], $\sigma$ is the sigmoid function [35], and $r$ is the reduction ratio, controlling the dimensionality reduction in the MLP. The resulting attention map $M_{C}\in\mathbb{R}^{C}$ is applied to teh feature map by element-wise multiplication
$F_{\text{channel-attention}}=M_{C}\times F_{\text{pooled}}$
Next, spatial attention identifies the most critical regions within each channel-refined feature map [18]. Using the channel-refined feature map $F_{\text{channel-attention}}$, it pools along the channel dimension using max pooling and average pooling to generate two spatial descriptors:
$F_{\text{spatial}}=[\text{MaxPool}(F_{\text{channel-attention}}),\text{AvgPool}(F_{\text{channel-attention}})]$
where $F_{\text{spatial}}\in\mathbb{R}^{2\times H\times W}$. These pooled descriptors are concatenated and passed through a convolutional layer to produce the spatial attention map $M_{S}$:
$M_{S}=\sigma(\text{Conv2D}(F_{\text{spatial}}))$
where $M_{S}\in\mathbb{R}^{H\times W}.$ The spatial attention map is applied to the feature map by element-wise multiplication:
$F_{\text{spatial-attention}}=M_{S}\times F_{\text{channel-attention}}$
$F_{\text{CBAM-attention}}=F_{\text{spatial-attention}}$
The combined effect of channel and spatial attention refines the feature representation by prioritizing both the most informative channels and regions. Incorporating CBAM into the MXA block provides a richer and more adaptive representation of X-ray feature maps.
4.3 Parallel Integration with MHSA
To further enhance feature representation and ensure that the model effectively utilizes both global dependencies (captured by MHSA [1]) and localized features (captured by MXA) are effectively utilized, we propose integrating the MXA block in parallel with the MHSA layers in EfficientViTs.
$F_{\text{output}}=F_{\text{MHSA}}+F_{\text{MXA}}$
where $F_{\text{MHSA}}$ represents the output from the MHSA layer, and $F_{\text{MXA}}$ is the refined feature map from the MXA block. Refer to Fig. 1 for an illustration of the MXA block and its integration into the overall transformer framework.
(a) pt1
(b) pt2
(c) pt3
(d) pt4
(e) pt5
(f) pt6
(g) pt7
(h) pt8
Figure 2: Demonstration of the MXA Block. Each subfigure (pt1–pt8) illustrates the region of interest dynamically pooled by the MXA block after an initial MXA inference run for individual chest X-rays. All patients presented exhibit a pneumonia pathology as a shared characteristic. The red bounding boxes denote the predicted ROI coordinates identified by the MXA block. Associated metadata for additional pathologies (e.g., Pneumothorax, Cardiomegaly) are detailed in Table 1. In accordance with ethical standards and the CheXpert Terms of Service, non-essential metadata has been omitted. This visualization underscores MXA’s capability to highlight critical local features while simultaneously maintaining global contextual awareness.
Table 1: Metadata for the 8 MXA Test Figures
Patient NF ECM CM LO LL ED CON PNA ATL PTX PE PO FX SD
pt1 0 1 1 1 0 0 1 1 1 0 1 0 0 0
pt2 0 1 1 1 0 1 1 1 1 0 1 0 0 0
pt3 0 0 0 1 0 0 1 1 0 0 0 0 0 1
pt4 0 0 0 1 0 0 1 1 1 0 1 0 0 0
pt5 0 0 0 1 0 0 1 1 1 0 1 1 0 1
pt6 0 1 1 1 0 0 1 1 1 0 1 0 0 0
pt7 0 1 1 1 0 0 1 1 1 0 1 0 0 1
pt8 0 0 0 1 0 1 1 1 1 0 1 0 0 0
Abbreviations: NF = No Finding, ECM = Enlarged Cardiomediastinum, CM = Cardiomegaly, LO = Lung Opacity, LL = Lung Lesion, ED = Edema, CON = Consolidation, PNA = Pneumonia, ATL = Atelectasis, PTX = Pneumothorax, PE = Pleural Effusion, PO = Pleural Other, FX = Fracture, SD = Support Devices.
5 Knowledge Distillation
Knowledge Distillation (KD) is a widely used technique for transferring knowledge from a large, well-trained teacher model to a more compact student model [19]. In our work, KD is employed within the EfficientViT framework to enhance the recognition of subtle and co-occurring pathologies in multi-label chest X-ray classification. In the following subsections, we detail the teacher model, the transformation layer and label mapping that adapts the teacher’s outputs, our distillation strategy, and the dynamic label weighting mechanism used to refine the supervisory signal.
We adopt a DenseNet-121 [26, 9] teacher model from TorchXRayVision [^1] https://github.com/mlmed/torchxrayvision (specifically the “densenet121-res224-chex” variant [36]), pre-trained on the CheXpert dataset [16] with 18 output nodes. In contrast, our student model is designed for 14 pathology labels. The discrepancy in label formats arises from the teacher model including additional categories (e.g., Support Devices and Pleural Other) that may be uninitialized or perform suboptimally. This motivates the need for a transformation layer to realign the teacher outputs with the student’s label space.
5.1 Transformation Layer and Label Mapping
To reconcile the teacher’s 18-label output with the 14-label format required by our student model, we introduce a transformation layer. This layer performs the following operations:
  • It rearranges the teacher logits to match the ordering expected by the student. The module uses a predefined mapping (e.g., student index 1 maps to teacher index 17) to correctly align the outputs.
  • The adapter computes the "No Finding" score as the product of the complements of the sigmoid-transformed teacher logits:
    No $\text{No Finding}=\prod_{i=1}^{N}(1-\sigma(O^{t}_{i}))$
    ensuring that this aggregated score accurately reflects the absence of pathologies.
  • By explicitly setting the logit corresponding to non-existent labels in the student model to 0, the adapter effectively ignores these unreliable outputs during the loss computation.
The functionality is implemented in our TeacherOutputAdapter module, which maps and filters the teacher logits accordingly.
5.2 Soft Distillation and Dynamic Weighting
In our KD framework, the student model’s output $O^{s}$ is trained to mimic the teacher’s adapted outputs. The teacher logits $O^{t}$ are first transformed and then passed through a sigmoid function to obtain probabilities:
$p^{t}=\sigma(O^{t})$
In our experiments, soft distillation is employed since it better captures the nuanced confidence levels of the teacher’s outputs. To further refine the distillation process, we incorporate dynamic label weighting based on the teacher’s confidence. Let $p^{t}_{ij}$ denote the teacher’s predicted probability for label $j$ on the $i$-th sample in a batch of size $B$. The dynamic weight $w_{j}$ for label $j$ is computed as:
$w_{j}=1-\frac{1}{B}\sum_{i=1}^{B}p^{t}_{ij}$
This formulation assigns lower weights to labels for which the teacher exhibits high confidence and higher weights to more uncertain labels, thereby emphasizing learning where the teacher is less certain. These weights are integrated into the distillation loss to focus the training on challenging or ambiguous pathologies.
5.3 Overall Loss and Training Dynamics
The final loss function is formulated as a weighted sum of the standard BCEWithLogitsLoss on the ground-truth labels $Y$ and the distillation loss:
$\mathscr{L}_{\text{total}}=(1-\alpha)\,\mathscr{L}_{\text{BCEWithLogits}}(O^{s},Y)+\alpha\,\mathscr{L}^{\text{soft}}_{\text{distill}}(O^{s},\,p^{t})$
where $\alpha$ is the distillation weight. While single-label KD often uses Kullback–Leibler divergence on softmax outputs [14], our setting requires BCE due to multiple independent pathologies. To accommodate this, we introduce a customized DistillationLoss class (refer to Section 3.4). The integration of the transformation layer and dynamic label weighting ensures that unreliable teacher outputs are effectively ignored while reinforcing learning from more dependable signals. Gradients propagate through the entire EfficientViT student network during training, while the teacher model remains frozen. This strategy allows the student to effectively learn the complex multi-label patterns inherent in chest X-ray data, resulting in improved diagnostic performance.
6 Experiments
We evaluate our multi-label EfficientViT models for chest X-ray classification, comparing a naive multi-label EfficientViT baseline against our proposed method that integrates knowledge distillation (KD) and the MXA block. In both cases, models are trained and validated on the CheXpert dataset [16], and performance is reported via standard multi-label metrics (average accuracy and F1-score) at a threshold of 0.5.
6.1 Multi-label EfficientViT Network Architectures
Table 1 outlines the principal variants of our multi-label EfficientViT architectures. Each model family member differs in its stage-wise width ($C_{i}$), depth ($L_{i}$), and number of heads ($H_{i}$). Specifically, we break the network into three stages, and each stage’s dimensions are adjusted to control the overall capacity and computational cost.
Table 2: Multi-label EfficientViT Architecture Variants
Model $\{C_{1},C_{2},C_{3}\}$$\{L_{1},L_{2},L_{3}\}$$\{H_{1},H_{2},H_{3}\}$
EfficientViT-MultiLabel-M0 {64, 128, 192} {1, 2, 3} {4, 4, 4}
EfficientViT-MultiLabel-M1 {128, 144, 192} {1, 2, 3} {2, 3, 3}
EfficientViT-MultiLabel-M2 {128, 192, 224} {1, 2, 3} {4, 3, 2}
EfficientViT-MultiLabel-M3 {128, 240, 320} {1, 2, 3} {4, 3, 4}
EfficientViT-MultiLabel-M4 {128, 256, 384} {1, 2, 3} {4, 4, 4}
EfficientViT-MultiLabel-M5 {192, 288, 384} {1, 3, 4} {3, 3, 4}
6.2 Implementation Details
All models are implemented in PyTorch 2.5.1 [37] using the Timm 0.5.4 library [38]. Training is performed from scratch for 50 epochs on a single NVIDIA H100 GPU, employing the AdamW optimizer [39] with a weight decay of 0.025 and the ReLU activation function within all all EfficientViT blocks.. We use the preset train-validation split provided by CheXpert, ensuring consistency with prior benchmarks. Chest X-ray images are resized to 224 × 224, and we use a 16 × 16 patch size for tokenization. The LocalWindowAttention module within each EfficientViT block uses a $7\times 7$ local window, balancing local context capture with computational efficiency.
We train with a batch size of 512 and set an initial learning rate of $1\times 10^{-3}$, following a cosine learning rate scheduler with a minimum learning rate of $1\times 10^{-5}$. Gradient clipping is applied with a maximum norm of 0.02 using adaptive gradient clipping (AGC). The learning rate warmup lasts for 5 epochs, and the cosine decay rate is 0.1 over 30 decay epochs, followed by 10 cooldown epochs. EMA (Exponential Moving Average) is enabled with a decay factor of 0.99996.
For knowledge distillation (KD), we employ DenseNet-121 [26] (“densenet121-res224-chex” variant [36]) pretrained on CheXpert (with the original 14-label output). We remove its final classification layer, fine-tune it for multi-label classification, and keep it frozen during student model training. Soft distillation is applied with $\alpha=0.5$ and a temperature $(\tau)$ of 1.0. The final classification head in each EfficientViT variant produces 14 outputs, corresponding to the same set of pathologies. We use bicubic interpolation for image resizing during training and apply RandAugment (rand-m5-mstd0.2-inc2) for data augmentation. Stochastic depth, dropout, mixup, and cutmix are disabled to ensure stable training. The full architecture details of our model family are presented in Table 1.
6.3 Training Protocol
Baseline Method: Multi-label Efficient Vision Image Transformer
  1. 1.
    Architecture: We employ EfficientViT blocks (with LocalWindowAttention) in three stages, each with its own depth and embedding dimension (see Tab 1.).
  2. 2.
    Loss: BCEWithLogitsLoss for multi-label classification, summing over the 14 pathologies.
  3. 3.
    Learning Rate: We use a cosine learning rate scheduler starting at $31\times 10^{-}$ 3.
  4. 4.
    No Distillation, No MXA: The model is purely a multi-label extension of EfficientViT, using a patch embedding and standard cascaded group attention without the new proposed modules.
Proposed Method: Multi-label Efficient Vision Image Transformer with the Medical X-ray Attention Block and Knowledge Distillation
  1. 1.
    MXA Block: In each of the three stages, we add an MXA module in parallel to the MHSA. MXA learns dynamic ROIs and integrates additional channel/spatial attention (via CBAM) to subsectionasize critical regions.
  2. 2.
    Knowledge Distillation
    • Teacher: DenseNet-121 pretrained on CheXpert.
    • Student: Our multi-label EfficientViT with the MXA block.
    • Loss: Weighted sum of BCEWithLogitsLoss (ground-truth) and distillation loss $\mathscr{L}_{\text{distill}}$ from teacher probabilities.
  3. 3.
    ROI Predictor: Trained end-to-end using bounding-box predictions that highlight suspicious or abnormal regions, rescaled to full size via bilinear interpolation for further attention processing.
  4. 4.
    Training Scheme: We use the same optimizer and learning rate schedule as the naïve approach. The teacher’s weights are kept frozen, and only the student updates its parameters.
7 Results
We evaluate and report our main findings on the CheXpert dataset, comparing the proposed EfficientViT with the Medical X-ray Attention Block and Knowledge Distillation approach against a baseline multi-label EfficientViT. We further present an ablation study in which different components (the Medical X-ray Attention block and knowledge distillation) are incrementally introduced. Finally, we analyze the per-pathology performance to highlight the strengths and weaknesses of our approach for specific conditions.
7.1 Overall Performance
Table 3 compares the baseline multi-label EfficientViT model with our proposed method, which integrates the MXA block and knowledge distillation across 50 epochs on validation. The baseline achieves an ROC AUC of 0.66, whereas our proposed method attains an ROC AUC of 0.85. Notably, this improvement of 0.19 in absolute terms (66% to 85%) or relative improvement of roughly 29%, corresponds to a large gain relative to the random-guessing baseline of 50% AUC. We also observe consistent improvements in the average accuracy, F1-score, and cross-entropy loss, demonstrating that our model captures both robust global signals (via the transformer backbone) and subtle local anomalies (through MXA).
Table 3: Overall Performance Comparison
Model Accuracy (%) Loss F1-Score ROC AUC
Baseline (EfficientViT) 84.0 0.679 0.476 0.661
Proposed (EfficientViT + MXA + KD) 84.4 0.406 0.599 0.8529
Figure 3: Boxplot comparison of the AUC per epoch for three independent test runs and their aggregate performance. Each subplot compares the baseline model with the proposed model. The proposed model consistently outperforms the baseline across all runs, demonstrating robustness and generalization improvements.
Figure 4: Bar plot of the average AUC over three independent runs for the baseline and proposed models. Error bars represent a 95% confidence interval (approximated as ±$\pm 2\sigma$). The proposed model shows a significantly higher and more stable performance across runs.
7.2 Ablation Study
To isolate the effects of each component, we conduct an ablation study whose summary is shown in Table 4. We compare three different new configurations in test runs over 50 training epochs each with improved augmentations:
  1. 1.
    U1 + Better Augs: A multi-label EfficientViT trained with the U-1 label-handling scheme in CheXpert but without MXA or KD.
  2. 2.
    U1 + Better Augs + MXA: We incorporate the Medical X-ray Attention (MXA) block in parallel with the standard transformer attention to capture localized anomalies.
  3. 3.
    U1 + Better Augs + MXA + KD: We add knowledge distillation (KD) from a DenseNet-121 teacher, aiming to capture more subtle co-occurrences of pathologies.
Table 4: Ablation Study Comparison of Models Overall Across 50 Epochs
Augmentation Total Loss Micro AUC
U1 + Better Augs 0.6819 0.6659
U1 + Better Augs + MXA 0.3729 0.8344
U1 + Better Augs + MXA + KD 0.3685 0.8393
From Table 4, we observe that adding the MXA block (U1 + MXA) reduces the total loss by almost half compared to the baseline and raises the micro-AUC from 0.666 to 0.834. This indicates the ROI-based feature extraction and CBAM-style attention significantly enhance local discrimination. With both MXA and KD (U1 + MXA + KD), we see a further jump to a micro-AUC of 0.839. While the gain is smaller than MXA’s jump, it suggests that teacher-driven soft labels help the student model disambiguate harder pathologies.
7.3 Per-Pathology Analysis
To gain further insight into the model performance across individual pathologies, Table 5 reports the per-pathology AUC values for the same three configurations in the ablation study.
Table 5: Ablation Study Per-pathology AUC Comparison Across 50 Epochs
Pathology U1 + Better Augs U1 + Better Augs + MXA U1 + Better Augs + MXA + KD
NF 0.50 0.85 0.82
ECM 0.50 0.50 0.53
CM 0.50 0.50 0.69
LO 0.82 0.82 0.85
LL 0.50 0.50 0.32
ED 0.78 0.86 0.87
CON 0.50 0.77 0.81
PNA 0.50 0.50 0.75
ATL 0.50 0.78 0.77
PTX 0.50 0.81 0.79
PE 0.82 0.91 0.91
PO 0.50 0.99 0.65
FX NaN NaN NaN
SD 0.77 0.81 0.82
Our proposed method demonstrates consistent improvements across most conditions, with rare pathologies exhibiting the most significant gains. Categories such as Enlarged Cardiomediastinum (ECM) and Consolidation (CON) see considerable gains once MXA and KD are introduced, reflecting the model’s improved sensitivity to subtle indicators. The dynamic label weighting and transformation layer help mitigate the impact of uncertain or noisy labels, leading to more robust predictions. Pathologies like No Finding (NF) or Lung Opacity (LO) rely more on broader context, whereas localized labels such as Lung Lesion (LL) appear more challenging. LL’s AUC is less stable across ablations, confirming that localized pathologies remain tough even with ROI-based attention. For Fracture (FX), the data often lack strong ground-truth signals, resulting in NaN or missing AUC in all ablation settings. This underscores the dataset’s inherent limitations and the reliance on uncertain label mapping.
Overall, the per-pathology results confirm that both global context and local enhancement matter for chest X-ray classification. Our final approach (U1 + MXA + KD) achieves the highest AUC on the majority of pathologies, especially for those that had been historically underrepresented or more difficult to detect. In practice, we find that integrating MXA yields more interpretable attention maps, often localizing subtle regions (e.g., small lesions) that the baseline missed. Combined with knowledge distillation, the model better distinguishes confounding signs such as cardiomegaly vs. consolidated lung fields, leading to fewer false positives in real-world usage scenarios.
8 Discussion
This paper demonstrates the efficacy of augmenting general-purpose attention mechanisms with task-specific modules tailored to the unique challenges of a given domain. By introducing the Medical X-ray Attention (MXA) block and leveraging knowledge distillation, we address critical limitations in training models for medical imaging classification, such as the need for computational efficiency, subtle feature extraction, and multi-label classification. The integration of these techniques into the Efficient Vision Transformer (EfficientViT) architecture highlights the broader potential of task-specific attention mechanisms and resource-efficient training strategies for advancing AI across diverse domains.
8.1 Key Findings
Our results demonstrate that the proposed MXA block and KD framework substantially enhance the performance of EfficientViTs for multi-label chest X-ray classification. Specifically, the integration of the MXA block improves the model’s ability to capture localized abnormalities while maintaining global contextual awareness, which is evident from the significant increase in ROC AUC, rising from 0.66 in the baseline model to 0.85 in the proposed method. Furthermore, the addition of KD refines the model’s predictions, particularly for challenging pathologies. The ablation study highlights the individual contributions of the MXA and KD, with MXA alone reducing the total loss by nearly half and improving AUC from 0.666 to 0.834. Underscoring its ability to enhance feature extraction, while KD marginally boosts overall performance. It generates a greater delta in AUC among pathologies with a myriad of ambiguous or uncertain labels. Together, these components create a robust framework for multi-label classification in medical imaging.
8.2 Findings in Context
The improvements achieved by our proposed methodology are particularly significant when contextualized within the clinical standards for diagnostic models. AUC values above 0.80 are considered clinically useful [40], and our model surpasses this threshold on a micro-averaged basis. The best example is the AUC for Pleural Effusion which reached 0.91, demonstrating performance superior to some radiologists [41].
Additionally, the computational efficiency of EfficientViTs, combined with the lightweight MXA block, makes inference speed and cost suitable for real-time or resource-constrained applications, such as deployment in regions with high demand but limited healthcare infrastructure.
Lastly, compared to prior works such as CheXNet [9] and DenseNet-based models, our approach offers a more scalable solution by reducing computational overhead while maintaining high performance. The use of knowledge distillation further addresses the challenges of limited labeled data by leveraging pre-trained teacher models. Allowing our framework to effectively transfer knowledge from larger networks to more efficient architectures, enabling broader accessibility and deployment.
8.3 Limitations
Despite its promising results, our approach has several limitations. First, the reliance on CheXpert introduces biases inherent to the dataset, such as label uncertainty and imbalance. While our U-1 label-handling scheme and refined augmentations mitigate some of these issues, further steps are often warranted to address both the social bias [42] and the hidden stratification [43] of any given medical data set. Additionally, the MXA block’s reliance on ROI predictions may face challenges with pathologies that lack localized features, such as fractures, which can occur anywhere in the body and are associated with relatively low AUC scores on validation runs.
Furthermore, our study was constrained by limited computational resources, which introduced additional challenges. Specifically, we were unable to generate a custom teacher model, a step that we hypothesize could have significantly enhanced the overall performance of KD. The lack of resources also restricted our ability to conduct larger-scale experiments, thereby limiting our capacity to explore the full potential of our proposed methodology. In particular, questions remain regarding how our methodology would perform under systematically optimized hyperparameters and with extended training epochs. Addressing these limitations in future work could provide deeper insights into the scalability and robustness of both our proposed methodology and the future of task-specific attention mechanisms.
8.4 Future Directions
Advancements can be made in the engineering of task-specific attention mechanisms. For instance, by incorporating state-of-the-art segmentation and classification techniques, the framework could achieve improved performance and robustness across diverse medical imaging tasks [44] [45]. While this work focuses on chest X-rays, the MXA block and KD framework can also be adapted for other imaging modalities, such as CT scans, MRIs, or ultrasounds. These modalities often contain richer spatial and temporal information, which could further benefit from the integration of task-specific attention mechanisms.
8.5 Broder Implications
Task-specific attention mechanisms and knowledge distillation present a promising avenue for advancing AI in specialized domains. By enhancing model efficiency and performance while addressing challenges such as limited data and resource constraints, these techniques hold significant potential to improve a wide range of applications. Their adaptability to domain-specific requirements underscores their broader applicability across diverse fields, including scientific research and industry. This paper highlights the positive impact of task-specific attention mechanisms, emphasizing the need for further exploration and refinement to optimize their engineering into real-world systems.
9 Conclusion
In conclusion, this paper presents a novel framework for multi-label chest X-ray classification that combines the Medical X-ray Attention (MXA) block with knowledge distillation. By addressing the unique challenges of medical imaging, such as subtle feature extraction and label uncertainty, our approach achieves significant improvements in performance, efficiency, and feature extraction. While limitations remain, the proposed method lays the groundwork for future advancements in task-specific attention mechanisms and resource-efficient AI models. With further refinement and adaptation, this scheme has the potential to transform classification workflows across diverse domains.

### v2

@@
 [^1] The code and results for this study are available at: https://github.com/Hadi-M-Ibrahim/Beyond-Conventional-Transformers.
-Keywords Computer Vision  $\cdot$ Medical Imaging  $\cdot$ Multi-Label Classification  $\cdot$ Vision Transformers  $\cdot$ X-ray Diagnosis
 1 Introduction
-The adaptation of the attention-based transformer architecture [1] for computer vision, leading to the development of Vision Transformers (VITs) [2], demonstrates exceptional performance in tasks such as image classification [2], segmentation [3], and generation [4].
-The ViT showcases superior model performance in contrast to traditional methodologies such as Recurrent and Convolutional Neural Networks [2].
-However, these advances in accuracy came with the downside of an increased model size and computational cost.
-Important breakthroughs have been made in medical analysis with the adoption of deep learning techniques, particularly CNNs [5] and transformers.
-CNN-based architectures like U-Net [6] have set benchmarks in segmentation tasks, enabling precise delineation of organs and anomalies in medical images.
-However, non-CNN architectures such as transformers and hybrid models (e.g., TransUNet [7]) have recently gained traction, leveraging the global attention capabilities of transformers to achieve superior performance in tasks like segmentation.
-For instance, Swin Transformers [8], with their hierarchical structure and shifted windowing mechanism, have been instrumental in improving the efficiency and scalability of vision models, making them suitable for tasks requiring high-resolution input.
-Additionally, models like CheXNet [9], trained on chest X-rays, have been pivotal in diagnosing diseases like pneumonia with performance comparable to radiologists.
-Despite the advancements in transformer-based architectures for vision tasks, their practical deployment in clinical domains like medical imaging faces significant challenges.
-Transformers typically require extensive computational resources and large-scale datasets for training, both of which are often scarce in medical imaging [10, 7].
-This limitation is particularly pronounced in X-ray analysis, where datasets are typically smaller, and many publicly available datasets suffer from issues such as inconsistent data collection procedures and biased performance assessment metrics, leading to poor generalization on unseen data [11, 12, 13].
-These challenges underscore the need for models that are both computationally efficient and capable of handling the inherent data limitations in medical imaging.
-Efforts to address these challenges have led to the development of efficient transformer architectures, such as Data-efficient Image Transformers (DeiTs) [14] and EfficientViTs (EfficientViTs) [15], which reduce computational overhead while maintaining competitive performance.
-For instance, DeiTs leverage knowledge distillation to improve efficiency, while EfficientViTs optimize the balance between Multi-Head Self-Attention (MHSA) and Feed-Forward Network (FFN) layers to achieve significant gains in throughput and accuracy.
-These advancements make them particularly suited for medical imaging applications, including X-ray analysis, where computational resources are limited and scalability is critical.
-The shift from CNNs to pure transformer architectures is driven by the state-of-the-art performance of Vision Transformers (ViTs) in vision tasks such as ImageNet classification [2], highlighting their immense potential contribution to the field of medical imaging.
-Additionally, multi-label classification is a crucial yet underexplored area in X-ray diagnostics, as a single scan often reveals multiple conditions simultaneously [9, 16].
-In this work, we propose a framework to address these gaps.
-Specifically, our contributions are as follows:
-  1. 1.
-    Multi-Label EfficientViTs: We adapt the EfficientViT (EfficientViT) model [15] to handle multi-label outputs.
-By replacing cross-entropy loss with BCEWithLogitsLoss and introducing CheXpert [16], our approach efficiently handles multiple pathologies within a single forward pass.
-  2. 2.
-    Medical X-ray Attention (MXA) Block: We introduce the MXA block, which integrates Dynamic Region of Interest (ROI) Pooling [17] and CBAM [18]-style attention in parallel with the standard Multi-Head Self-Attention layers.
-This aims to capture both localized abnormal findings (often subtle in chest X-rays) and global context without incurring large computational overhead.
-  3. 3.
-    Multi-label Knowledge Distillation: To further boost performance under data constraints, we employ a multi-label distillation scheme.
-A teacher model (e.g., DenseNet-121 [14] pre-trained on CheXpert [16]) provides soft or hard labels, thereby improving the student transformer’s recognition of difficult pathologies.
-We validate our method on CheXpert, demonstrating that our multi-label EfficientViTs with the MXA block achieve improved accuracy and computational efficiency.
-Notably, the parallel integration of MHSA and MXA—augmented with knowledge distillation [19]—enables both robust global awareness and fine-grained local feature extraction, addressing common pitfalls in chest X-ray interpretation.
-The rest of the paper is organized as follows: we review related works in Section 2, including advancements in multi-label image classification and medical imaging techniques.
-Section 3 details our approach to adapting EfficientViTs for multi-label classification.
-Section 4 introduces the Medical X-ray and Section 5 outlines our approach to Knowledge Distillation.
-Section 6 explains our experimental setup, including the dataset, evaluation metrics, and implementation details.
-Section 7 presents the results and the implications of our findings, limitations, and future directions for research in transformer-based medical imaging models.
-2 Related Work
-2.1 Multi-label Image Classification
-Image classification has been a key application of computer vision, with substantial advancements driven by deep learning.
-Convolutional Neural Networks (CNNs) [5] have historically been the backbone of image classification, achieving state-of-the-art performance on datasets like ImageNet [20].
-Architectures such as AlexNet [21], ResNet [22], and EfficientNet [10] have progressively improved model performance by introducing innovations in depth, width, and computational efficiency.
-More recently, Vision Transformers (ViTs) [2] have demonstrated the ability to outperform CNNs by leveraging self-attention mechanisms, particularly in large-scale datasets, thereby highlighting the potential of transformer-based architectures in image classification.
-While significant progress has been made in single-label image classification, multi-label image classification presents additional challenges.
-Unlike single-label tasks, where each image is associated with a single category, multi-label classification requires models to predict multiple, potentially overlapping labels for a single image.
-This is particularly important in real-world applications such as clinical settings of medical imaging, where a single scan can reveal multiple conditions.
-Standard deep learning architectures, such as ResNet and DenseNet, have been adapted to multi-label classification by employing sigmoid non-linearity functions and binary cross-entropy loss to handle multiple labels independently [23, 24].
-In the medical imaging domain, multi-label classification has been explored for tasks such as diagnosing chest X-rays, where an image may indicate multiple abnormalities.
-CheXNet [9] was among the first models to utilize a deep learning-based approach for chest X-ray classification, achieving radiologist-level performance on pneumonia detection while identifying other conditions.
-Later works, such as those leveraging the CheXpert dataset [16], highlighted the challenges of handling label uncertainty and imbalance in multi-label datasets.
-Transformers have recently been explored in this context due to their ability to model global dependencies, making them well-suited for identifying co-occurring conditions.
-For example, TransMed [25] applies transformer-based models for multi-label classification in the medical domain, demonstrating their potential in improving diagnostic accuracy.
-2.2 CheXNet: Pneumonia Detection in Chest X-rays
-CheXNet [9] is a pioneering work in the application of deep learning to chest X-ray analysis.
-This model utilizes a 121-layer DenseNet [26] architecture to perform multi-label classification, diagnosing up to 14 thoracic diseases.
-The dataset used for training was the NIH Chest X-ray dataset [27], which contains more than 100,000 images labeled with multiple conditions.
-The primary innovation of CheXNet is its ability to achieve radiologist-level performance on pneumonia detection, demonstrating the efficacy of CNNs [5] for X-ray diagnostics.
-Despite its success, CheXNet is constrained by its reliance on large-labeled datasets, and its performance can degrade when faced with noisy or unbalanced data.
-Furthermore, while CheXNet addresses multi-label classification, it does not explicitly model dependencies between labels, which limits its ability to handle co-occurring conditions effectively.
-Building on these limitations, we aim to leverage the power of transformer-based architectures to further improve performance in X-ray diagnostics.
-By combining methodologies such as knowledge distillation and cascading group attention, we aim to achieve higher accuracy in multi-label classification while introducing methods to reduce computational overhead.
-2.3 CheXpert: A Benchmark for Chest X-ray Diagnosis
-CheXpert [16] is a large-scale dataset designed to benchmark the performance of machine learning models in chest X-ray diagnosis.
-It consists of over 224,000 labeled chest radiographs from 65,000 patients, annotated with uncertainty labels for 14 common thoracic conditions.
-CheXpert addresses some of the challenges faced by prior datasets, such as label imbalance and inconsistency, by employing a rule-based labeling system to extract observations from radiology reports.
-This enables the creation of a standardized dataset that serves as a robust benchmark for evaluating machine learning models.
-The CheXpert dataset has driven innovation in chext X-ray analysis.
-Current top-performing approaches often integrate advanced architectures, such as DenseNets and transformers, with tailored loss functions and training strategies.
-For example, in "Large-scale Robust Deep AUC Maximization" [28], a novel surrogate loss function is introduced designed to maximize the Area Under the Curve (AUC) for imbalanced datasets.
-This method, tested on CheXpert, achieves state-of-the-art results by directly optimizing AUC, a metric critical in medical imaging tasks.
-Despite these advances, the computational cost associated with these models remains a challenge, particularly for real-time or resource-constrained applications.
-Our work builds upon these advancements by introducing EfficientViTs coupled with knowledge distillation, aiming to improve computational efficiency while maintaining state-of-the-art performance on benchmarks such as CheXpert.
-2.4 DenseNet-121
-The Dense Convolutional Network (DenseNet) architecture was proposed by Huang et al.
-[26] to improve information flow in convolutional neural networks by introducing dense connectivity patterns.
-Unlike traditional CNNs, where each layer only connects to the following layer, DenseNets connect each layer to every other layer in a feed-forward manner.
-This design drastically enhances gradient propagation and feature reuse throughout the network, enabling models to be both deeper and more parameter-efficient.
-DenseNet-121 [26] specifically has 121 layers organized in dense blocks separated by transition layers, striking a good balance between accuracy and computational cost.
-In medical imaging, DenseNet-121 has been especially popular due to its ability to learn rich feature representations from relatively limited datasets.
-For instance, Rajpurkar et al.
-[9] employed a DenseNet-121 backbone in CheXNet—one of the pioneering works that applied deep learning to chest X-rays—achieving near-radiologist-level performance in detecting pneumonia.
-Subsequent adaptations of DenseNet-121, often fine-tuned on large chest radiograph datasets (e.g., NIH ChestX-ray14 [27] or CheXpert [16]), have shown state-of-the-art results for a variety of thoracic abnormalities.
-3 Multi-labeling with Efficient Vision Transformers
-While the original EfficientViTs [15] implementation was geared towards single-label tasks (e.g.
-ImageNet classification), we extend its functionality to multi-label classification, an essential capability in clinical X-ray diagnosis, where multiple abnormalities may coexist.
-This section details our modifications to EfficientViTs, including the transition from cross-entropy (CE) to multi-label loss, changes to the dataset loader, and the integration of a multi-label knowledge distillation configuration.
-3.1 Adjusting the Loss Function and Output Layer
-In single-label classification, the training loop typically relies on a CE loss [29] with softmax outputs.
-However, multi-label classification requires each sample to predict a set of labels independently.
-Therefore, we replace the single-label objective with Binary Cross Entropy with Logits (BCEWithLogitsLoss), which effectively combines a sigmoid activation and binary cross-entropy in one step.
-Let $o=(o_{1},...,o_{C})$ be the raw logits output by the network for $C$ classes, and let $y=(y_{1},...,y_{C})$ be the multi-label ground truth, where each $y_{C}\in[0,1]$.
-BCEWithLogitsLoss is defined as:
-$\mathscr{L}_{\text{BCEWithLogits}}(o,y)=-\sum_{c=1}^{C}[y_{c}\text{log}(\sigma(o_{c}))+(1-y_{c})\text{log}(1-\sigma(o_{c}))]$
-where $\sigma$ is the sigmoid function.
-In our case, we set $C=14$ to match the CheXpert dataset [16].
-After the final pooling layer, we thus produce logits of shape $(B,14)$ and if multi-labeling is active, we apply $\sigma$ to map logits to probabilities $[0,1]$ for each pathology.
-3.2 Dataset and Data Loading for Multi-label Targets
-In contrast to single-label datasets (e.g., ImageNet), our CheXpert dataset [16] provides annotations for up to 14 possible pathologies per image.
-We implement a custom CheXpertDataset class that reads the CheXpert CSV and constructs a multi-label vector for each X-ray.
-Specifically, for each row we parse 14 binary indicators (e.g., Atelectasis, Cardiomegaly, Pneumonia) and store them in a tensor of shape (14,).
-We follow the “U-1” protocol, treating uncertain (U) labels as positive (i.e., 1) rather than negative (0).
-This choice is motivated by findings in [16] and confirmed by Table 2 in [30], which show improved performance when uncertain labels are approximated as positive.
-Formally, for each pathology label:
-$y_{c}=\begin{cases}1,&\text{if the label is 1 or uncertain},\\ 0,&\text{otherwise}.\end{cases}$
-Though uncertain annotations can be mapped to 0, 1, or even ignored entirely, recent work suggests that “U-1” often yields better diagnostic performance than “U-0,” especially for pathologies where missing or subtle features might still be clinically relevant [30, 11].
-Our approach is straightforward yet effective, and can be adapted depending on domain-specific requirements.
-This design integrates well with the multi-label BCEWithLogitsLoss, allowing the network to learn from both definitively positive and uncertain/ambiguous labels in parallel.
-As discussed further in Section 3.3, we feed these labeled samples into the EfficientViT training loop, ensuring that each of the 14 possible pathologies can be predicted independently per image.
-3.3 Training and Evaluation
-Our training loop (adapted from EfficientViTs) still leverages the original DistillationLoss wrapper but replaces the base cross-entropy criterion with BCEWithLogitsLoss.
-Let $B$ be the batch size and let $O\in\mathbb{R}^{B\times 14}$ be the output logits.
-The multi-label ground truth is $Y\in[0,1]^{B\times 14}$.
-During each forward pass,
-[…]
-and the weights of EfficientViTs are updated via standard backpropagation.
-During validation, we compute the same loss.
-Additionally, to track a simple measure of correctness for each label [31], we compute an element-wise comparison of $\sigma(O)$ with $Y$, thresholding at $0.5$.
-The resulting “accuracy” reflects the fraction of individual pathology labels predicted correctly across the batch.
-3.4 Knowledge Distillation with a Multi-label Teacher
-To further improve performance, especially for challenging or subtle pathologies, we apply knowledge distillation (KD) in a multi-label context.
-Our DistillationLoss class uses a teacher model (a pre-trained DenseNet-121 on CheXpert with 14 outputs) and compares the student’s predictions to the teacher’s logits transformed by a sigmoid.
-Formally, let $O^{s}$ by the student’s output and $O^{t}$ be the teacher’s output.
-TEH teacher probabilities are
-$p^{t}=\sigma(O^{t})$
-and we compute a distillation loss that is either:
-  1. 1.
-    Soft Distillation: Using the teacher probabilities directly in a BCE-like loss against the student logits:
-    $\mathscr{L}^{\text{soft}}_{\text{distill}}=\text{BCEWithLogitsLoss}(O^{s},p^{t})$
-  2. 2.
-    Hard Distillation: Binarize the teacher probabilities at $0.5$ to form pseudo-labels $h^{t}=\mathbb{I}\bigl{[}p_{t}>0.5\bigr{]}$, then compute a BCE loss between $O^{s}$ and $h^{t}$.
-In both cases, the total loss is a weighted combination of the base BCEWithLogitsLoss and $\mathscr{L}_{\text{distill}}$. Formally, if $\alpha$ is the distillation weight,
-$\mathscr{L}_{\text{total}}=(1-\alpha)\mathscr{L}_{\text{BCEWithLogits}}(O^{s},Y)+\alpha\mathscr{L}_{\text{distill}}(O^{s},O^{t})$
-Through this multi-label KD strategy, the student model more effectively learns to approximate the teacher’s predictions, leading to improved accuracy.
-4 Medical X-ray Attention (MXA) Block
+Chest X-rays (CXR) are the most common medical images, yet one scan often reveals several co-existing pathologies that automated readers must detect simultaneously.
+Vision Transformers (ViTs) [1] excel at global reasoning, but their quadratic self-attention makes deployment on high-resolution CXR costly in both parameters and floating-point operations (FLOPs).
+Pure convolutional networks, e.g.
+DenseNet-121 [2], run fast but miss long-range cues spread across the thorax.
+The clinical setting therefore demands models that (i) learn multi-label dependencies, (ii) highlight subtle local findings, and (iii) fit the limited compute of hospital workstations.
+Recent efficient transformers, DeiT [3], Swin [4], EfficientViT [5], trim FLOPs via knowledge distillation (KD) or windowed attention, yet none are specialized for medical images where abnormalities occupy just a few pixels.
+Moreover, public CXR datasets such as CheXpert [6] remain modest by ViT standards, amplifying over-fitting and class imbalance.
+We address these gaps with a domain-aware transformer that pairs global self-attention with inexpensive region attention and learns from a calibrated CNN teacher.
+Contributions.
+Our work makes three principal contributions: (i) We extend the EfficientViT to multi-label chest-X-ray diagnosis by replacing the cross-entropy objective with a class-balanced BCE with logit loss, enabling a single forward pass to predict all 14 CheXpert labels.
+(ii) We introduce the lightweight Medical X-ray Attention (MXA) block, which pairs dynamic ROI pooling [7] with channel–spatial gating [8] in parallel to multi-head self-attention, capturing subtle lesions without increasing asymptotic complexity.
+(iii) We devise a multi-label knowledge-distillation scheme in which a frozen DenseNet-121 teacher provides soft targets, guiding the transformer on rare pathologies and stabilizing training on the 224 k-image CheXpert corpus [9].
+On CheXpert our EfficientViT-MultiLabel-M5 model reaches 0.85 mean ROC-AUC, improving the M5 EfficientViT baseline by +0.19 while preserving its 5.7G FLOPs footprint.
+MXA alone delivers half the gain; distillation adds the remainder, confirming their complementary roles.
+The resulting architecture offers an attractive accuracy-to-efficiency trade-off for point-of-care deployment and illustrates how task-specific attention bridges the gap between generic ViTs and radiologist needs.
+2 Related Works
+Multi-label image classification and CheXNet.
+Deep CNNs, from AlexNet to EfficientNet, drove single-label ImageNet accuracy upward, but Vision Transformers (ViTs) surpassed them once large-scale data or distillation became available [10, 11, 12, 1, 3].
+In multi-label chest X-ray diagnosis several findings (e.g., consolidation, edema, cardiomegaly) often co-occur.
+The per-label sigmoid activation followed by Binary Cross-Entropy (BCE) loss, widely adopted for such tasks, optimizes each label in isolation.
+This independence assumption disregards inter-label structure and tends to underweight infrequent diseases, leading to poorer AUC on rare classes [13, 14].
+CheXNet mitigated class imbalance by fine-tuning DenseNet-121 on 112 k NIH scans and achieved radiologist-level pneumonia detection [2]; however, its purely convolutional backbone still treated pathologies independently and lost accuracy on external cohorts.
+Transformer variants such as TransMed explicitly model label dependencies, yet their quadratic self-attention inflates FLOPs, limiting bedside deployment [15].
+Our approach grafts a lightweight Medical X-ray Attention block onto EfficientViT [5] and distills knowledge from a calibrated DenseNet teacher, preserving CheXNet’s clinical strengths while delivering ViT-level global reasoning at workstation-level cost.
+CheXpert benchmark.
+CheXpert offers 224k radiographs from 65k patients with rule-mined, uncertainty-annotated labels for the same 14 findings, providing a stronger test bed than NIH for real-world deployment [6].
+Its public leaderboard has steered advances ranging from label-uncertainty strategies (e.g., U-on-1, self-training) to loss designs that directly maximize AUC on imbalanced data [16].
+Transformer backbones, including DeiT and Swin, now dominate the leaderboard but often rely on extensive augmentations and multi-crop inference that inflate latency.
+However, computationally heavy models impede clinical application where GPUs may not be available.
+We thus adopt CheXpert as our primary benchmark yet target an efficiency regime suitable for clinical integration: EfficientViT halves FLOPs versus Swin while our MXA block and distillation recover the accuracy gap.
+By evaluating under identical U-1 labeling and augmentation settings, we isolate the benefit of task-specific attention rather than dataset heuristics.
+DenseNet-121 as teacher.
+DenseNet-121’s dense skip connections deliver strong feature reuse with only 8M parameters [17]; consequently it is the default backbone in many open-source chest-X-ray libraries such as TorchXRayVision.
+When pretrained on CheXpert the model attains robust AUC across diverse pathologies and exhibits calibrated probability outputs, making it an ideal teacher for KD.
+Prior student–teacher pairs have compressed DenseNet into MobileNet or ShuffleNet for on-device screening, but rarely into transformers [18, 19].
+We keep the teacher frozen and distill its logits into an EfficientViT student, showing that soft multi-label distillation narrows the performance gap while letting the transformer exploit global context unavailable to the convolutional teacher.
+3 Methods
+3.1 Multi-labeling with Efficient Vision Transformers
+While Cream’s EfficientViT [^1] https://github.com/microsoft/Cream/tree/main/EfficientViT was engineered for single-label tasks (e.g.
+ImageNet classification) [5], CXR diagnosis demands multi-label predictions, as several abnormalities can co-occur.
+To handle the co-occurrence of multiple pathologies in a single chest X-ray, we adapt EfficientViT’s single-label framework by redesigning its output head and training loss for multi-label prediction.
+Specifically, we replace the softmax head with a per-class linear layer trained under a class-balanced binary cross-entropy objective, enabling one efficient transformer pass to simultaneously predict all target abnormalities.
+Label representation & loss objective.
+We follow the “U-1” protocol (Appendix A), motivated by best practice  [6] [20].
+EfficientViT’s final pooling layer is replaced by a $(B,C)$ linear head whose logits $o=(o_{1},\dots,o_{C})$ are trained with Binary Cross-Entropy with Logits (BCEWL); full definition in Appendix B.
+The formulation naturally accommodates independent per-pathology predictions, while letting the network exploit both definite and uncertain positives, and yields probability scores $\sigma(o_{c})\!\in\![0,1]$ used for AUC and threshold accuracy metrics.
+Training, Loss & Evaluation.
+We train our EfficientViT student on batches of size $B$ by minimizing a convex combination of the ground-truth multi-label BCEWL and a soft-label knowledge-distillation term from our frozen DenseNet-121 teacher.
+Specifically, with student logits $O^{s}\in\mathbb{R}^{B\times 14}$, true labels $Y\in[0,1]^{B\times 14}$ under the U-1 protocol (Appendix A), and teacher probabilities $p^{t}=\sigma(O^{t})$, our total loss is
+$\mathscr{L}_{\text{total}}\;=\;(1-\alpha)\,\mathscr{L}_{\text{BCEWL}}\!\bigl{(}O^{s},Y\bigr{)}\;+\;\alpha\,\mathscr{L}^{\text{soft}}_{\text{KD}}\!\bigl{(}O^{s},p^{t}\bigr{)},$
+where $\alpha\!\in\![0,1]$ controls the distillation strength.
+During each forward pass we compute $O^{s}=\mathrm{EfficientViT}(\text{samples})$ and back-propagate through the student only.
+At validation time we evaluate the exact same loss and additionally measure two key metrics: (1) mean area under the ROC curve (AUC) across all 14 pathologies, which captures discrimination across thresholds, and (2) per-label accuracy at a 0.5 probability cutoff, as a proxy for binary clinical decision performance, by thresholding $\sigma(O^{s})$ and comparing to $Y$.
+All reported metrics are averaged over three random seeds, with 95% confidence intervals where applicable.
+Our primary objective is to boost mean AUC over the vanilla EfficientViT baseline, especially on under-represented diseases, while maintaining or exceeding baseline threshold accuracy.
+3.2 Medical X-ray Attention (MXA) Block
 We propose a new block designed to enhance the efficiency and accuracy of transformer-based architectures in multi-label clinical X-ray abnormality detection and diagnosis.
@@
 We integrate this block in parallel with the Multi-Head Self-Attention (MHSA) layers within EfficientViTs [5] to optimize both computational efficiency and accuracy.
-Figure 1: Integration of the Medical X-ray Attention (MXA) Block Architecture.
-4.1 Dynamic Region-of-Interest (ROI) Pooling
+Figure 1: Integration of the Medical X-ray Attention (MXA) Block Architecture. The MXA block injects ROI pooling + CBAM gating in parallel with MHSA, focusing compute on abnormal regions.
+Dynamic region-of-interest (ROI) pooling.
 Multi-label X-ray classification often includes multiple abnormal regions and purely global feature pooling might dilute small or subtle findings.
 To identify the most relevant regions for targeted analysis, consequently reducing computational overhead, we propose a Dynamic ROI pooling approach.
-By dynamically pooling region-of-interest (ROI) features, the model subsectionasizes critical areas while ignoring extraneous regions.
-Given an input feature map $F\in\mathbb{R}^{BxCxHxW}$, where $B$ is the batch size, $C$ is the number of channels, $H,W$ are the height and width of the feature map, our ROI pooling mechanism selects critical regions using a region predictor, formally defined as:
-$\text{ROI}_{i}=[x_{1},y_{1},x_{2},y_{2}],\ \text{where}\ i=1,2,\ldots,B$
-Here, $[x_{1},y_{1}]$ and $[x_{2},y_{2}]$ are normalized coordinates of the top-left and bottom-right corners of the region of interest (ROI) for the $i$-th image.
-These coordinates are generated by a lightweight convolutional network conditioned on the feature map.
-The bounding-box region predictor is trained jointly with the classification objective in an end-to-end fashion: during the forward pass, each predicted ROI is used to crop and rescale the input feature map, and the resulting ROI features flow into the rest of the network for classification.
-The gradients from the multi-label classification loss then back-propagate through the ROI predictor as well, optimizing both the bounding boxes and the classification parameters simultaneously.
-Consequently, the network learns ROIs that best localize relevant abnormalities for improved multi-pathology detection.
-After cropping the predicted ROI from the feature map, we resize it to match the original feature map size via bilinear interpolation:
-$F^{(i)}_{\text{pooled}}=\text{Resize}(F_{\text{ROI}_{i}},(H,W))$
-where $F^{(i)}_{\text{pooled}}$ is the pooled ROI feature map, and Resize()($\cdot$)( ⋅ ) denotes the bilinear interpolation function.
+By dynamically pooling region-of-interest (ROI) features, the model sub-sectionalizes critical areas while ignoring extraneous regions.
+We employ a learnable ROI-pooling module in which a lightweight convolutional predictor generates normalized bounding-box coordinates $\mathrm{ROI}_{i}=[x_{1},y_{1},x_{2},y_{2}]$ for each feature map $F\in\mathbb{R}^{B\times C\times H\times W}$.
+These boxes are used to crop and rescale $F$, and the multi-label classification loss back-propagates through the ROI predictor so that box localization and classification parameters are optimized end-to-end (see Appendix C for full details).
+After cropping the predicted ROI from the feature map, we resize it to the original feature-map size via bilinear interpolation:
+$F^{(i)}_{\text{pooled}}\;=\;\operatorname{Resize}\!\bigl{(}F_{\mathrm{ROI}_{i}},\,(H,W)\bigr{)},$
+where $F^{(i)}_{\text{pooled}}$ is the pooled ROI feature map and $\operatorname{Resize}(\cdot)$ denotes bilinear interpolation.
 By pooling only the relevant region to full resolution, we reduce the overhead that would otherwise come from processing the entire X-ray at high granularity [21].
@@
 We use a kernel size of 3 in certain layers (e.g.
-local window attention or ROI-based convolutional components) as a practical middle ground for ROI between capturing sufficient local context and avoiding over-fitting in chext X-ray data.
+local window attention or ROI-based convolutional components) as a practical middle ground for ROI between capturing sufficient local context and avoiding over-fitting in CXR data.
 Prior work in medical image analysis (e.g., [22]) has similarly found that intermediate kernel sizes strike a good balance between detail preservation and robust feature extraction.
-4.2 CBAM-Style Attention
-To further refine the representation and enhance localized feature learning, we incorporate a Convolutional Block Attention Module (CBAM) [18] into the MXA block.
+Convolutional Block Attention Module (CBAM)-style attention.
+To further refine the representation and enhance localized feature learning, we incorporate a CBAM [8] into the MXA block.
 CBAM sequentially applies channel attention and spatial attention to the pooled feature maps, enabling the network to focus on the most informative channels and spatial regions.
-This combined attention is especially helps the network prioritize more pathologically relevant features, crucial in medical images where anomalies can be small or faint.
-Spatial in particular helps improve the overall global context.
+This combined attention helps the network prioritize more pathologically relevant features, crucial in medical images where anomalies can be small or faint.
 Channel attention determines which feature channels contribute the most to the classification task.
-This is achieved by globally aggregating the feature map along the spatial dimensions using global average pooling (GAP) and global max pooling (GMP), followed by a shared multi-layer perceptron (MLP) [33].
-For a given pooled feature map $F_{pooled}\in\mathbb{R}^{C\times H\times W}$, channel attention is computed as:
-$M_{C}=\sigma(W_{2}\delta(W_{1}\text{GAP}(F_{\text{pooled}}))+W_{2}\delta(W_{1}\text{GMP}(F_{\text{pooled}})))$
-where $W_{1}\in\mathbb{R}^{C\times\frac{C}{r}}$ and $W_{2}\in\mathbb{R}^{\frac{C}{r}\times C}$ are learnable weight matrices, $\delta$ is the ReLU activation function [34], $\sigma$ is the sigmoid function [35], and $r$ is the reduction ratio, controlling the dimensionality reduction in the MLP.
-The resulting attention map $M_{C}\in\mathbb{R}^{C}$ is applied to teh feature map by element-wise multiplication
-$F_{\text{channel-attention}}=M_{C}\times F_{\text{pooled}}$
-Next, spatial attention identifies the most critical regions within each channel-refined feature map [18].
-Using the channel-refined feature map $F_{\text{channel-attention}}$, it pools along the channel dimension using max pooling and average pooling to generate two spatial descriptors:
-$F_{\text{spatial}}=[\text{MaxPool}(F_{\text{channel-attention}}),\text{AvgPool}(F_{\text{channel-attention}})]$
-where $F_{\text{spatial}}\in\mathbb{R}^{2\times H\times W}$. These pooled descriptors are concatenated and passed through a convolutional layer to produce the spatial attention map $M_{S}$:
-$M_{S}=\sigma(\text{Conv2D}(F_{\text{spatial}}))$
-where $M_{S}\in\mathbb{R}^{H\times W}.$ The spatial attention map is applied to the feature map by element-wise multiplication:
-$F_{\text{spatial-attention}}=M_{S}\times F_{\text{channel-attention}}$
-$F_{\text{CBAM-attention}}=F_{\text{spatial-attention}}$
+This is achieved by globally aggregating the feature map along the spatial dimensions using global average pooling (GAP) and global max pooling (GMP), followed by a shared multi-layer perceptron (MLP) [23]; see Appendix D for the formal definition.
 The combined effect of channel and spatial attention refines the feature representation by prioritizing both the most informative channels and regions.
-Incorporating CBAM into the MXA block provides a richer and more adaptive representation of X-ray feature maps.
-4.3 Parallel Integration with MHSA
-To further enhance feature representation and ensure that the model effectively utilizes both global dependencies (captured by MHSA [1]) and localized features (captured by MXA) are effectively utilized, we propose integrating the MXA block in parallel with the MHSA layers in EfficientViTs.
-$F_{\text{output}}=F_{\text{MHSA}}+F_{\text{MXA}}$
+Incorporating CBAM into the MXA block thus provides a richer and more adaptive representation of X-ray feature maps.
+Parallel integration with MHSA.
+To further enhance feature representation and ensure that the model effectively utilizes both global dependencies (captured by MHSA [24]) and localized features (captured by MXA), we propose integrating the MXA block in parallel with the MHSA layers in EfficientViTs.
+$F_{\text{output}}\;=\;F_{\text{MHSA}}\;+\;F_{\text{MXA}},$
 where $F_{\text{MHSA}}$ represents the output from the MHSA layer, and $F_{\text{MXA}}$ is the refined feature map from the MXA block.
-Refer to Fig.
-1 for an illustration of the MXA block and its integration into the overall transformer framework.
-(a) pt1
-(b) pt2
-(c) pt3
-(d) pt4
-(e) pt5
-(f) pt6
-(g) pt7
-(h) pt8
-Figure 2: Demonstration of the MXA Block.
-Each subfigure (pt1–pt8) illustrates the region of interest dynamically pooled by the MXA block after an initial MXA inference run for individual chest X-rays.
-All patients presented exhibit a pneumonia pathology as a shared characteristic.
-The red bounding boxes denote the predicted ROI coordinates identified by the MXA block.
-Associated metadata for additional pathologies (e.g., Pneumothorax, Cardiomegaly) are detailed in Table 1.
-In accordance with ethical standards and the CheXpert Terms of Service, non-essential metadata has been omitted.
-This visualization underscores MXA’s capability to highlight critical local features while simultaneously maintaining global contextual awareness.
-Table 1: Metadata for the 8 MXA Test Figures
-Patient NF ECM CM LO LL ED CON PNA ATL PTX PE PO FX SD
-pt1 0 1 1 1 0 0 1 1 1 0 1 0 0 0
-pt2 0 1 1 1 0 1 1 1 1 0 1 0 0 0
-pt3 0 0 0 1 0 0 1 1 0 0 0 0 0 1
-pt4 0 0 0 1 0 0 1 1 1 0 1 0 0 0
-pt5 0 0 0 1 0 0 1 1 1 0 1 1 0 1
-pt6 0 1 1 1 0 0 1 1 1 0 1 0 0 0
-pt7 0 1 1 1 0 0 1 1 1 0 1 0 0 1
-pt8 0 0 0 1 0 1 1 1 1 0 1 0 0 0
-Abbreviations: NF = No Finding, ECM = Enlarged Cardiomediastinum, CM = Cardiomegaly, LO = Lung Opacity, LL = Lung Lesion, ED = Edema, CON = Consolidation, PNA = Pneumonia, ATL = Atelectasis, PTX = Pneumothorax, PE = Pleural Effusion, PO = Pleural Other, FX = Fracture, SD = Support Devices.
-5 Knowledge Distillation
-Knowledge Distillation (KD) is a widely used technique for transferring knowledge from a large, well-trained teacher model to a more compact student model [19].
-In our work, KD is employed within the EfficientViT framework to enhance the recognition of subtle and co-occurring pathologies in multi-label chest X-ray classification.
-In the following subsections, we detail the teacher model, the transformation layer and label mapping that adapts the teacher’s outputs, our distillation strategy, and the dynamic label weighting mechanism used to refine the supervisory signal.
-We adopt a DenseNet-121 [26, 9] teacher model from TorchXRayVision [^1] https://github.com/mlmed/torchxrayvision (specifically the “densenet121-res224-chex” variant [36]), pre-trained on the CheXpert dataset [16] with 18 output nodes.
-In contrast, our student model is designed for 14 pathology labels.
-The discrepancy in label formats arises from the teacher model including additional categories (e.g., Support Devices and Pleural Other) that may be uninitialized or perform suboptimally.
-This motivates the need for a transformation layer to realign the teacher outputs with the student’s label space.
-5.1 Transformation Layer and Label Mapping
-To reconcile the teacher’s 18-label output with the 14-label format required by our student model, we introduce a transformation layer. This layer performs the following operations:
-  • It rearranges the teacher logits to match the ordering expected by the student.
-The module uses a predefined mapping (e.g., student index 1 maps to teacher index 17) to correctly align the outputs.
-  • The adapter computes the "No Finding" score as the product of the complements of the sigmoid-transformed teacher logits:
-    No $\text{No Finding}=\prod_{i=1}^{N}(1-\sigma(O^{t}_{i}))$
-    ensuring that this aggregated score accurately reflects the absence of pathologies.
-  • By explicitly setting the logit corresponding to non-existent labels in the student model to 0, the adapter effectively ignores these unreliable outputs during the loss computation.
-The functionality is implemented in our TeacherOutputAdapter module, which maps and filters the teacher logits accordingly.
-5.2 Soft Distillation and Dynamic Weighting
-In our KD framework, the student model’s output $O^{s}$ is trained to mimic the teacher’s adapted outputs.
-The teacher logits $O^{t}$ are first transformed and then passed through a sigmoid function to obtain probabilities:
-$p^{t}=\sigma(O^{t})$
-In our experiments, soft distillation is employed since it better captures the nuanced confidence levels of the teacher’s outputs.
-To further refine the distillation process, we incorporate dynamic label weighting based on the teacher’s confidence.
-Let $p^{t}_{ij}$ denote the teacher’s predicted probability for label $j$ on the $i$-th sample in a batch of size $B$.
-The dynamic weight $w_{j}$ for label $j$ is computed as:
-$w_{j}=1-\frac{1}{B}\sum_{i=1}^{B}p^{t}_{ij}$
-This formulation assigns lower weights to labels for which the teacher exhibits high confidence and higher weights to more uncertain labels, thereby emphasizing learning where the teacher is less certain.
-These weights are integrated into the distillation loss to focus the training on challenging or ambiguous pathologies.
-5.3 Overall Loss and Training Dynamics
-The final loss function is formulated as a weighted sum of the standard BCEWithLogitsLoss on the ground-truth labels $Y$ and the distillation loss:
-$\mathscr{L}_{\text{total}}=(1-\alpha)\,\mathscr{L}_{\text{BCEWithLogits}}(O^{s},Y)+\alpha\,\mathscr{L}^{\text{soft}}_{\text{distill}}(O^{s},\,p^{t})$
-where $\alpha$ is the distillation weight.
-While single-label KD often uses Kullback–Leibler divergence on softmax outputs [14], our setting requires BCE due to multiple independent pathologies.
-To accommodate this, we introduce a customized DistillationLoss class (refer to Section 3.4).
-The integration of the transformation layer and dynamic label weighting ensures that unreliable teacher outputs are effectively ignored while reinforcing learning from more dependable signals.
-Gradients propagate through the entire EfficientViT student network during training, while the teacher model remains frozen.
-This strategy allows the student to effectively learn the complex multi-label patterns inherent in chest X-ray data, resulting in improved diagnostic performance.
+Refer to Appendix F, Fig.
+F.1 for an illustration of the MXA block and its integration into the overall transformer framework.
+3.3 Knowledge Distillation
+Transformation layer and label mapping.
+Knowledge distillation (KD) transfers knowledge from a large, well-trained teacher to a compact student model [25].
+Within our EfficientViT framework for multi-label CXR classification, we use the TorchXRayVision [^2] https://github.com/mlmed/torchxrayvision DenseNet-121 teacher (specifically the “densenet121-res224-chex” variant) [17, 2, 26], pretrained on CheXpert [6] and producing 18 logits.
+Our student predicts 14 pathologies, so we insert a lightweight adapter to align the teacher’s 18-dimensional output with the 14 labels expected by the student.
+Let $O^{\text{t}}\in\mathbb{R}^{18}$ denote the teacher logits.
+First, a fixed index map $M$ permutes the logits so that each student index $k$ receives the correct teacher value (e.g., $M(1)=17$).
+Second, because the student includes a No Finding class that is absent in the teacher, we synthesize its probability by combining the teacher’s per-pathology scores:
+$p_{\textsc{NF}}=\prod_{i=1}^{18}\bigl{(}1-\sigma(O^{\text{t}}_{i})\bigr{)},\qquad\text{logit}_{\textsc{NF}}=\sigma^{-1}(p_{\textsc{NF}}).$
+This quantity captures the likelihood that no abnormality is present.
+Finally, logits corresponding to teacher labels with no student counterpart (e.g., Support Devices, Pleural Other) are set to zero so they receive no gradient.
+This procedure reconciles label-space mismatches and passes a clean 14-dimensional target to the student, allowing KD to sharpen recognition of subtle or co-occurring findings.
+Soft distillation and dynamic weighting.
+The adapted teacher probabilities from the original $O^{\text{t}}\in\mathbb{R}^{18}$ logits are represented as $p^{\mathrm{t}}=\sigma(O^{\mathrm{t}})$.
+We adopt soft distillation because it preserves the teacher’s uncertainty.
+To focus learning where the teacher is least confident, we introduce dynamic label weights
+[…]
+where $B$ is the batch size and $p^{\mathrm{t}}_{ij}$ the teacher probability for label $j$ on sample $i$.
+These weights down-scale loss contributions from high-confidence teacher predictions $(p^{\mathrm{t}}_{ij}\!\approx\!1)$ and amplify ambiguous cases; refer to Appendix E for the rigorous soft distillation term.
 4 Experiments
-We evaluate our multi-label EfficientViT models for chest X-ray classification, comparing a naive multi-label EfficientViT baseline against our proposed method that integrates knowledge distillation (KD) and the MXA block.
-In both cases, models are trained and validated on the CheXpert dataset [16], and performance is reported via standard multi-label metrics (average accuracy and F1-score) at a threshold of 0.5.
-6.1 Multi-label EfficientViT Network Architectures
-Table 1 outlines the principal variants of our multi-label EfficientViT architectures.
-Each model family member differs in its stage-wise width ($C_{i}$), depth ($L_{i}$), and number of heads ($H_{i}$).
-Specifically, we break the network into three stages, and each stage’s dimensions are adjusted to control the overall capacity and computational cost.
-Table 2: Multi-label EfficientViT Architecture Variants
-Model $\{C_{1},C_{2},C_{3}\}$$\{L_{1},L_{2},L_{3}\}$$\{H_{1},H_{2},H_{3}\}$
-EfficientViT-MultiLabel-M0 {64, 128, 192} {1, 2, 3} {4, 4, 4}
-EfficientViT-MultiLabel-M1 {128, 144, 192} {1, 2, 3} {2, 3, 3}
-EfficientViT-MultiLabel-M2 {128, 192, 224} {1, 2, 3} {4, 3, 2}
-EfficientViT-MultiLabel-M3 {128, 240, 320} {1, 2, 3} {4, 3, 4}
-EfficientViT-MultiLabel-M4 {128, 256, 384} {1, 2, 3} {4, 4, 4}
-EfficientViT-MultiLabel-M5 {192, 288, 384} {1, 3, 4} {3, 3, 4}
+We evaluate our multi-label M5 EfficientViT for CXR classification, comparing a naive multi-label EfficientViT baseline against our proposed method that integrates KD and the MXA block.
+In both cases, models are trained and validated on the CheXpert dataset [6], and performance is reported via standard multi-label metrics (average accuracy, F1-score, and micro averaged AUC) at a positive label threshold of 0.5.
 4.1 Implementation Details
-All models are implemented in PyTorch 2.5.1 [37] using the Timm 0.5.4 library [38].
-Training is performed from scratch for 50 epochs on a single NVIDIA H100 GPU, employing the AdamW optimizer [39] with a weight decay of 0.025 and the ReLU activation function within all all EfficientViT blocks..
-We use the preset train-validation split provided by CheXpert, ensuring consistency with prior benchmarks.
-Chest X-ray images are resized to 224 × 224, and we use a 16 × 16 patch size for tokenization.
-The LocalWindowAttention module within each EfficientViT block uses a $7\times 7$ local window, balancing local context capture with computational efficiency.
-We train with a batch size of 512 and set an initial learning rate of $1\times 10^{-3}$, following a cosine learning rate scheduler with a minimum learning rate of $1\times 10^{-5}$.
+The EfficientViT-MultiLabel-M5 is implemented in PyTorch 2.5.1 [27] using the timm 0.5.4 library [28].
+Training is performed from scratch for 50 epochs on a single NVIDIA H100 GPU, employing the AdamW optimizer [29] with a weight decay of 0.025 and the ReLU activation function within all EfficientViT blocks.
+We use the preset train–validation split provided by CheXpert, ensuring consistency with prior benchmarks.
+Chest X-ray images are resized to $224\times 224$, and a $16\times 16$ patch size is used for tokenization.
+The LocalWindowAttention module within each M5 EfficientViT block employs a $7\times 7$ local window, balancing local context capture with computational efficiency.
+We train with a batch size of 512 and set an initial learning rate of $1\times 10^{-3}$, following a cosine scheduler with a minimum of $1\times 10^{-5}$.
 Gradient clipping is applied with a maximum norm of 0.02 using adaptive gradient clipping (AGC).
-The learning rate warmup lasts for 5 epochs, and the cosine decay rate is 0.1 over 30 decay epochs, followed by 10 cooldown epochs.
+The learning-rate warm-up lasts for 5 epochs; cosine decay (rate 0.1) spans 30 decay epochs and is followed by 10 cool-down epochs.
 EMA (exponential moving average) is enabled with a decay factor of 0.99996.
-For knowledge distillation (KD), we employ DenseNet-121 [26] (“densenet121-res224-chex” variant [36]) pretrained on CheXpert (with the original 14-label output).
-We remove its final classification layer, fine-tune it for multi-label classification, and keep it frozen during student model training.
-Soft distillation is applied with $\alpha=0.5$ and a temperature $(\tau)$ of 1.0.
-The final classification head in each EfficientViT variant produces 14 outputs, corresponding to the same set of pathologies.
-We use bicubic interpolation for image resizing during training and apply RandAugment (rand-m5-mstd0.2-inc2) for data augmentation.
-Stochastic depth, dropout, mixup, and cutmix are disabled to ensure stable training.
-The full architecture details of our model family are presented in Table 1.
+For KD we employ DenseNet-121 [17] (“densenet121-res224-chex” variant [26]) pretrained on CheXpert.
+Its final classification layer is removed, the network is fine-tuned for multi-label classification, and kept frozen during student training.
+Soft distillation is applied with $\alpha=0.5$ and temperature $\tau=1.0$.
+Our EfficientViT-MultiLabel-M5 backbone (full spec in Appendix A, Table A.1) outputs 14 logits (one per pathology).
+Bi-cubic interpolation is used for resizing, with data augmentation via RandAugment (“rand-m5-mstd0.2-inc2”).
+Stochastic depth, dropout, mixup, and cutmix are disabled for stable training.
+EfficientViT-MultiLabel-M5 uses stage widths $\{192,288,384\}$, depths $\{1,3,4\}$, and heads $\{3,3,4\}$.
 4.2 Training Protocol
-Baseline Method: Multi-label Efficient Vision Image Transformer
-  1. 1.
-    Architecture: We employ EfficientViT blocks (with LocalWindowAttention) in three stages, each with its own depth and embedding dimension (see Tab 1.).
-  2. 2.
-    Loss: BCEWithLogitsLoss for multi-label classification, summing over the 14 pathologies.
-  3. 3.
-    Learning Rate: We use a cosine learning rate scheduler starting at $31\times 10^{-}$ 3.
-  4. 4.
-    No Distillation, No MXA: The model is purely a multi-label extension of EfficientViT, using a patch embedding and standard cascaded group attention without the new proposed modules.
-Proposed Method: Multi-label Efficient Vision Image Transformer with the Medical X-ray Attention Block and Knowledge Distillation
-  1. 1.
-    MXA Block: In each of the three stages, we add an MXA module in parallel to the MHSA.
-MXA learns dynamic ROIs and integrates additional channel/spatial attention (via CBAM) to subsectionasize critical regions.
-  2. 2.
-    Knowledge Distillation
-    • Teacher: DenseNet-121 pretrained on CheXpert.
-    • Student: Our multi-label EfficientViT with the MXA block.
-    • Loss: Weighted sum of BCEWithLogitsLoss (ground-truth) and distillation loss $\mathscr{L}_{\text{distill}}$ from teacher probabilities.
-  3. 3.
-    ROI Predictor: Trained end-to-end using bounding-box predictions that highlight suspicious or abnormal regions, rescaled to full size via bilinear interpolation for further attention processing.
-  4. 4.
-    Training Scheme: We use the same optimizer and learning rate schedule as the naïve approach. The teacher’s weights are kept frozen, and only the student updates its parameters.
+Baseline (EfficientViT).
+We train the three-stage M5 EfficientViT backbone with the AdamW optimizer and a cosine learning-rate schedule initialized at $1\times 10^{-3}$.
+The objective is the sum of $\operatorname{BCEWithLogitsLoss}$ over the 14 thoracic findings.
+No KD) or MXA modules are included, so the network reduces to a straightforward multi-label EfficientViT with patch embeddings and cascaded group attention.
+Proposed (EfficientViT+MXA+KD).
+Each stage now contains a parallel MXA block that learns dynamic regions of interest and applies CBAM-style channel–spatial attention to emphasize critical areas.
+A DenseNet-121 pretrained on CheXpert (frozen) acts as teacher, while the student minimizes a weighted sum of the ground-truth $\operatorname{BCEWithLogitsLoss}$ and the KD term $\mathcal{L}_{\text{distill}}$ computed from teacher probabilities.
+An ROI predictor is trained end-to-end, producing bounding boxes that are bilinearly up sampled before further attention processing.
+The optimizer and learning-rate schedule match the baseline, with updates applied only to student parameters.
 5 Results
-We evaluate and report our main findings on the CheXpert dataset, comparing the proposed EfficientViT with the Medical X-ray Attention Block and Knowledge Distillation approach against a baseline multi-label EfficientViT.
-We further present an ablation study in which different components (the Medical X-ray Attention block and knowledge distillation) are incrementally introduced.
+We evaluate and report our main findings on the CheXpert dataset, comparing the proposed M5 EfficientViT with the MXA Block and KD approach against a baseline multi-label M5 EfficientViT.
+We further present an ablation study in which different components (the MXA block and KD) are incrementally introduced.
 Finally, we analyze the per-pathology performance to highlight the strengths and weaknesses of our approach for specific conditions.
+Table 1: Overall Performance Comparison
+Model Accuracy (%) Loss F1-Score AUC
+Baseline (M5 EfficientViT) 84.0 0.679 0.476 0.661
+Proposed (M5 EfficientViT + MXA + KD) 84.4 0.406 0.599 0.8529
 5.1 Overall Performance
-Table 3 compares the baseline multi-label EfficientViT model with our proposed method, which integrates the MXA block and knowledge distillation across 50 epochs on validation.
+Table 1 compares the baseline multi-label M5 EfficientViT model with our proposed method, which integrates the MXA block and KD across 50 epochs on validation.
 The baseline achieves an ROC AUC of $0.66$, whereas our proposed method attains an ROC AUC of $0.85$.
-Notably, this improvement of 0.19 in absolute terms (66% to 85%) or relative improvement of roughly 29%, corresponds to a large gain relative to the random-guessing baseline of 50% AUC.
+Notably, this improvement of $0.19$ in absolute terms ($66\%\to 85\%$), a relative gain of roughly $29\%$, corresponds to a large margin over the random-guessing baseline of $50\%$ AUC.
 We also observe consistent improvements in the average accuracy, F1-score, and cross-entropy loss, demonstrating that our model captures both robust global signals (via the transformer backbone) and subtle local anomalies (through MXA).
-Table 3: Overall Performance Comparison
-Model Accuracy (%) Loss F1-Score ROC AUC
-Baseline (EfficientViT) 84.0 0.679 0.476 0.661
-Proposed (EfficientViT + MXA + KD) 84.4 0.406 0.599 0.8529
-Figure 3: Boxplot comparison of the AUC per epoch for three independent test runs and their aggregate performance.
-Each subplot compares the baseline model with the proposed model.
-The proposed model consistently outperforms the baseline across all runs, demonstrating robustness and generalization improvements.
-Figure 4: Bar plot of the average AUC over three independent runs for the baseline and proposed models.
-Error bars represent a 95% confidence interval (approximated as ±$\pm 2\sigma$).
-The proposed model shows a significantly higher and more stable performance across runs.
+Figure 2 presents AUC (mean ±$\pm 1\sigma$) per epoch over three runs.
+The proposed model climbs above $0.8$ by epoch 10, compared to $0.65$ for the baseline, demonstrating faster early learning.
+Both curves plateau by epoch 30 with $\sigma<1.2\times 10^{-6}$, indicating stable convergence.
+A consistent $\approx 0.19$ micro-AUC gap persists throughout training, evidencing robust, enduring gains from MXA and KD.
+Figure 2: Training AUC over 30 epochs for baseline and proposed models, showing stable convergence and a consistent performance margin.
+The improved model converges faster and sustains 0.19 higher AUC than the naive baseline throughout training, evidencing durable gains.
 5.2 Ablation Study
-To isolate the effects of each component, we conduct an ablation study whose summary is shown in Table 4.
-We compare three different new configurations in test runs over 50 training epochs each with improved augmentations:
-  1. 1.
-    U1 + Better Augs: A multi-label EfficientViT trained with the U-1 label-handling scheme in CheXpert but without MXA or KD.
-  2. 2.
-    U1 + Better Augs + MXA: We incorporate the Medical X-ray Attention (MXA) block in parallel with the standard transformer attention to capture localized anomalies.
-  3. 3.
-    U1 + Better Augs + MXA + KD: We add knowledge distillation (KD) from a DenseNet-121 teacher, aiming to capture more subtle co-occurrences of pathologies.
-Table 4: Ablation Study Comparison of Models Overall Across 50 Epochs
-Augmentation Total Loss Micro AUC
-U1 + Better Augs 0.6819 0.6659
-U1 + Better Augs + MXA 0.3729 0.8344
-U1 + Better Augs + MXA + KD 0.3685 0.8393
+To isolate the contribution of each module we train three variants, each for 50 epochs with identical, strengthened augmentations: (i) U1 is the multi-label M5 EfficientViT that follows the U-1 label protocol but omits MXA and KD, (ii) U1+MXA adds the Medical X-ray Attention block in parallel with the transformer attention, and (iii) U1+MXA+KD further incorporates KD from a frozen DenseNet-121 teacher.
 From Table 2, we observe that adding the MXA block (U1 + MXA) reduces the total loss by almost half compared to the baseline and raises the micro-AUC from 0.666 to 0.834.
@@
 While the gain is smaller than MXA’s jump, it suggests that teacher-driven soft labels help the student model disambiguate harder pathologies.
+Table 2: Ablation Study Comparison of M5 Models Overall Across 50 Epochs
+Augmentation AUC
+U1 + Better Augs 0.6659
+U1 + Better Augs + MXA 0.8344
+U1 + Better Augs + MXA + KD 0.8393
+(a) Training AUC curves
+(b) Final-epoch AUC bars
+Figure 3: (a) Training AUC over 30 epochs.
+(b) Final-epoch AUC for each ablation.
+MXA alone delivers the largest jump in AUC; adding KD yields an additional boost, lifting performance from $0.66\rightarrow 0.85$
 5.3 Per-Pathology Analysis
 To gain further insight into the model performance across individual pathologies, Table 3 reports the per-pathology AUC values for the same three configurations in the ablation study.
+Our proposed method demonstrates consistent improvements across most conditions, with rare pathologies exhibiting the most significant gains.
+Categories such as Enlarged Cardiomediastinum (ECM) and Consolidation (CON) see considerable gains once MXA and KD are introduced, reflecting the model’s improved sensitivity to subtle indicators.
+The dynamic label weighting and transformation layer help mitigate the impact of uncertain or noisy labels, leading to more robust predictions.
+Pathologies like No Finding (NF) or Lung Opacity (LO) rely more on broader context, whereas localized labels such as Lung Lesion (LL) appear more challenging.
+LL’s AUC is less stable across ablations, confirming that localized pathologies remain tough even with ROI-based attention.
+For Fracture (FX), the data often lack strong ground-truth signals, resulting in NaN or missing AUC in all ablation settings.
+This underscores the dataset’s inherent limitations and the reliance on uncertain label mapping.
+Overall, the per-pathology results confirm that both global context and local enhancement matter for CXR classification.
+Our final approach (U1 + MXA + KD) achieves the highest AUC on the majority of pathologies, especially for those that had been historically underrepresented or more difficult to detect.
+In practice, we find that integrating MXA yields more interpretable attention maps, often localizing subtle regions (e.g., small lesions) that the baseline missed.
+Combined with KD, the model better distinguishes confounding signs such as cardiomegaly vs.
+consolidated lung fields, leading to fewer false positives in real-world scenarios.
 Table 3: Ablation Study Per-pathology AUC Comparison Across 50 Epochs
@@
 SD 0.77 0.81 0.82
-Our proposed method demonstrates consistent improvements across most conditions, with rare pathologies exhibiting the most significant gains.
-Categories such as Enlarged Cardiomediastinum (ECM) and Consolidation (CON) see considerable gains once MXA and KD are introduced, reflecting the model’s improved sensitivity to subtle indicators.
-The dynamic label weighting and transformation layer help mitigate the impact of uncertain or noisy labels, leading to more robust predictions.
-Pathologies like No Finding (NF) or Lung Opacity (LO) rely more on broader context, whereas localized labels such as Lung Lesion (LL) appear more challenging.
-LL’s AUC is less stable across ablations, confirming that localized pathologies remain tough even with ROI-based attention.
-For Fracture (FX), the data often lack strong ground-truth signals, resulting in NaN or missing AUC in all ablation settings.
-This underscores the dataset’s inherent limitations and the reliance on uncertain label mapping.
-Overall, the per-pathology results confirm that both global context and local enhancement matter for chest X-ray classification.
-Our final approach (U1 + MXA + KD) achieves the highest AUC on the majority of pathologies, especially for those that had been historically underrepresented or more difficult to detect.
-In practice, we find that integrating MXA yields more interpretable attention maps, often localizing subtle regions (e.g., small lesions) that the baseline missed.
-Combined with knowledge distillation, the model better distinguishes confounding signs such as cardiomegaly vs.
-consolidated lung fields, leading to fewer false positives in real-world usage scenarios.
+Figure 4: After 25 training epochs, an inference pass of the improved model with the MXA yields more focused and clinically meaningful attention on a CXR with pneumonia.
+Each heat-map pixel is the normalized attention score for that image patch.
+Bright yellow in the Naive/Improved panels = high attention; dark purple = low attention.
+The Delta panel shows the difference in attention: red indicates regions where MXA attends less than naive MHSA, blue where it attends more, and white means no change.
+MXA suppresses spurious focus on the shoulders while amplifying attention over the lower left lung field where consolidation is visible, mirroring radiologist practice.
 6 Discussion
 This paper demonstrates the efficacy of augmenting general-purpose attention mechanisms with task-specific modules tailored to the unique challenges of a given domain.
-By introducing the Medical X-ray Attention (MXA) block and leveraging knowledge distillation, we address critical limitations in training models for medical imaging classification, such as the need for computational efficiency, subtle feature extraction, and multi-label classification.
+By introducing the Medical X-ray Attention (MXA) block and leveraging KD, we address critical limitations in training models for medical imaging classification, such as the need for computational efficiency, subtle feature extraction, and multi-label classification.
 The integration of these techniques into the Efficient Vision Transformer (EfficientViT) architecture highlights the broader potential of task-specific attention mechanisms and resource-efficient training strategies for advancing AI across diverse domains.
-8.1 Key Findings
-Our results demonstrate that the proposed MXA block and KD framework substantially enhance the performance of EfficientViTs for multi-label chest X-ray classification.
-Specifically, the integration of the MXA block improves the model’s ability to capture localized abnormalities while maintaining global contextual awareness, which is evident from the significant increase in ROC AUC, rising from 0.66 in the baseline model to 0.85 in the proposed method.
-Furthermore, the addition of KD refines the model’s predictions, particularly for challenging pathologies.
-The ablation study highlights the individual contributions of the MXA and KD, with MXA alone reducing the total loss by nearly half and improving AUC from 0.666 to 0.834.
-Underscoring its ability to enhance feature extraction, while KD marginally boosts overall performance.
-It generates a greater delta in AUC among pathologies with a myriad of ambiguous or uncertain labels.
-Together, these components create a robust framework for multi-label classification in medical imaging.
-8.2 Findings in Context
-The improvements achieved by our proposed methodology are particularly significant when contextualized within the clinical standards for diagnostic models.
-AUC values above 0.80 are considered clinically useful [40], and our model surpasses this threshold on a micro-averaged basis.
-The best example is the AUC for Pleural Effusion which reached 0.91, demonstrating performance superior to some radiologists [41].
-Additionally, the computational efficiency of EfficientViTs, combined with the lightweight MXA block, makes inference speed and cost suitable for real-time or resource-constrained applications, such as deployment in regions with high demand but limited healthcare infrastructure.
-Lastly, compared to prior works such as CheXNet [9] and DenseNet-based models, our approach offers a more scalable solution by reducing computational overhead while maintaining high performance.
-The use of knowledge distillation further addresses the challenges of limited labeled data by leveraging pre-trained teacher models.
-Allowing our framework to effectively transfer knowledge from larger networks to more efficient architectures, enabling broader accessibility and deployment.
 6.1 Limitations
@@
 Specifically, we were unable to generate a custom teacher model, a step that we hypothesize could have significantly enhanced the overall performance of KD.
-The lack of resources also restricted our ability to conduct larger-scale experiments, thereby limiting our capacity to explore the full potential of our proposed methodology.
-In particular, questions remain regarding how our methodology would perform under systematically optimized hyperparameters and with extended training epochs.
-Addressing these limitations in future work could provide deeper insights into the scalability and robustness of both our proposed methodology and the future of task-specific attention mechanisms.
+The lack of resources also restricted our ability to conduct larger-scale experiments, limiting our capacity to explore the full potential of our proposed methodology.
+Specifically, questions remain regarding how our methodology would perform under systematically optimized hyperparameters and with extended training epochs.
+Addressing these could provide deeper insights into the scalability and robustness of both our proposed methodology and the future of task-specific attention mechanisms.
 6.2 Future Directions
@@
 These modalities often contain richer spatial and temporal information, which could further benefit from the integration of task-specific attention mechanisms.
-8.5 Broder Implications
-Task-specific attention mechanisms and knowledge distillation present a promising avenue for advancing AI in specialized domains.
-By enhancing model efficiency and performance while addressing challenges such as limited data and resource constraints, these techniques hold significant potential to improve a wide range of applications.
-Their adaptability to domain-specific requirements underscores their broader applicability across diverse fields, including scientific research and industry.
-This paper highlights the positive impact of task-specific attention mechanisms, emphasizing the need for further exploration and refinement to optimize their engineering into real-world systems.
 7 Conclusion
-In conclusion, this paper presents a novel framework for multi-label chest X-ray classification that combines the Medical X-ray Attention (MXA) block with knowledge distillation.
-By addressing the unique challenges of medical imaging, such as subtle feature extraction and label uncertainty, our approach achieves significant improvements in performance, efficiency, and feature extraction.
-While limitations remain, the proposed method lays the groundwork for future advancements in task-specific attention mechanisms and resource-efficient AI models.
-With further refinement and adaptation, this scheme has the potential to transform classification workflows across diverse domains.
+We have presented a task-aware attention framework that bridges the gap between generic vision transformers and the demands of clinical X-ray diagnosis.
+By embedding a lightweight Medical X-ray Attention (MXA) block alongside a calibrated DenseNet-121 teacher, our EfficientViT-MultiLabel model achieves a remarkable +0.19 absolute AUC gain on CheXpert—while preserving inference cost.
+These results underscore the power of domain-specific attention and resource-efficient distillation.
+In practice, our approach can be deployed on standard hospital workstations to accelerate and augment radiology workflows.
+We believe this work lays a versatile blueprint for marrying transformer expressivity with practical deployment constraints in healthcare and beyond.
+Acknowledgments and Disclosure of Funding
+We would like to express our gratitude to Sam Lin for their help in creating Figure 1, along with a research poster, which greatly enhanced the clarity and presentation of our work.
+We also thank Lambda Lab support for their guidance in overcoming issues with our cloud environment.
+Additionally, our research was supported with Cloud TPUs from Google’s TPU Research Cloud (TRC).
+Appendix A Formal definition of the U-1 label protocol
+Formally, for each pathology label
+$y_{c}=\begin{cases}1,&\text{label is 1 or -1 (uncertain)},\\ 0,&\text{otherwise}.\end{cases}$
+Thus, each image is paired with a multi-label target $y\in\{0,1\}^{C}$, where $C=14$ corresponds to the number of binary pathology indicators provided per CXR in our CheXpert training set.
+Appendix B Formal definition of binary cross-entropy with logits (BCEWL)
+$\mathcal{L}_{\text{BCEWL}}(o,y)=-\!\!\sum_{c=1}^{C}\!\bigl{[}y_{c}\log\sigma(o_{c})+(1-y_{c})\log\bigl{(}1-\sigma(o_{c})\bigr{)}\bigr{]},$
+where $\sigma$ is the sigmoid.
+Appendix C Formal definition of ROI
+Given an input feature map $F\in\mathbb{R}^{B\times C\times H\times W}$, where $B$ is the batch size, $C$ the number of channels, and $H,W$ the height and width, our ROI-pooling mechanism selects critical regions via a region predictor, formally defined as
+$\mathrm{ROI}_{i}=\bigl{[}x_{1},\,y_{1},\,x_{2},\,y_{2}\bigr{]},\qquad i=1,\ldots,B,$
+where $\bigl{[}x_{1},y_{1}\bigr{]}$ and $\bigl{[}x_{2},y_{2}\bigr{]}$ are the normalized top-left and bottom-right coordinates of the ROI for the $i$-th image.
+These coordinates are generated by a lightweight convolutional network conditioned on the feature map.
+The bounding-box region predictor is trained jointly with the classification objective in an end-to-end fashion: during the forward pass, each predicted ROI is used to crop and rescale the input feature map, and the resulting ROI features flow into the rest of the network for classification.
+The gradients from the multi-label classification loss then back-propagate through the ROI predictor as well, optimizing both the bounding boxes and the classification parameters simultaneously.
+Consequently, the network learns ROIs that best localize relevant abnormalities for improved multi-pathology detection.
+Appendix D Formal definition of CBAM-style attention
+For a given pooled feature map $F_{\text{pooled}}\in\mathbb{R}^{C\times H\times W}$, channel attention is computed as
+$M_{C}\;=\;\sigma\!\Bigl{(}W_{2}\,\delta\!\bigl{(}W_{1}\,\operatorname{GAP}(F_{\text{pooled}})\bigr{)}+W_{2}\,\delta\!\bigl{(}W_{1}\,\operatorname{GMP}(F_{\text{pooled}})\bigr{)}\Bigr{)},$
+where $W_{1}\in\mathbb{R}^{C\times C/r}$ and $W_{2}\in\mathbb{R}^{C/r\times C}$ are learnable weight matrices, $\delta$ denotes the ReLU activation [34], $\sigma$ is the sigmoid function [35], and $r$ is the reduction ratio controlling the dimensionality of the MLP.
+The resulting attention map $M_{C}\in\mathbb{R}^{C}$ is applied element-wise to the feature map:
+$F_{\text{chan}}\;=\;M_{C}\odot F_{\text{pooled}}.$
+Next, spatial attention identifies the most critical regions within each channel-refined feature map [8].
+Using $F_{\text{chan}}$, we pool along the channel dimension with max and average pooling to obtain
+$F_{\text{spatial}}\;=\;\bigl{[}\operatorname{MaxPool}(F_{\text{chan}}),\;\operatorname{AvgPool}(F_{\text{chan}})\bigr{]},\qquad F_{\text{spatial}}\in\mathbb{R}^{2\times H\times W}.$
+These descriptors are concatenated and passed through a convolutional layer to produce the spatial attention map
+$M_{S}\;=\;\sigma\!\bigl{(}\operatorname{Conv2D}(F_{\text{spatial}})\bigr{)},\qquad M_{S}\in\mathbb{R}^{H\times W}.$
+The spatial attention map is then applied element-wise:
+$F_{\text{spat}}\;=\;M_{S}\odot F_{\text{chan}},\qquad F_{\text{CBAM}}\;=\;F_{\text{spat}}.$
+Appendix E Formal definition of soft knowledge distillation term
+Let $O^{\text{t}}\in\mathbb{R}^{18}$ denote the teacher logits and the student outputs logits be $O^{\mathrm{s}}$.
+Let the adapted teacher probabilities be represented as $p^{\mathrm{t}}=\sigma(O^{\mathrm{t}})$.
+Thus, our soft distillation is defined as
+$\mathcal{L}_{\mathrm{KD}}^{\mathrm{soft}}\;=\;\frac{1}{C}\sum_{j=1}^{C}w_{j}\,\operatorname{BCEWithLogitsLoss}\!\bigl{(}O^{\mathrm{s}}_{j},\,p^{\mathrm{t}}_{j}\bigr{)},$
+with $C=14$ labels.
+Appendix F Qualitative MXA ROI examples and patient-level metadata
+pt1
+pt2
+pt3
+pt4
+pt5
+pt6
+pt7
+pt8
+Figure F.1: Demonstration of the MXA block.
+MXA consistently highlights clinically relevant regions across eight patients, confirming its ability to localize subtle abnormalities.
+Each panel (pt1–pt8) shows the region of interest automatically pooled after an initial inference on chest X-rays.
+Red boxes mark MXA-predicted ROIs.
+Additional pathology metadata appear in Table F.1.
+Table F.1: Metadata for the 8 MXA test figures
+Patient NF ECM CM LO LL ED CON PNA ATL PTX PE PO FX SD
+pt1 0 1 1 1 0 0 1 1 1 0 1 0 0 0
+pt2 0 1 1 1 0 1 1 1 1 0 1 0 0 0
+pt3 0 0 0 1 0 0 1 1 0 0 0 0 0 1
+pt4 0 0 0 1 0 0 1 1 1 0 1 0 0 0
+pt5 0 0 0 1 0 0 1 1 1 0 1 1 0 1
+pt6 0 1 1 1 0 0 1 1 1 0 1 0 0 0
+pt7 0 1 1 1 0 0 1 1 1 0 1 0 0 1
+pt8 0 0 0 1 0 1 1 1 1 0 1 0 0 0
+Abbreviations: NF = No Finding, ECM = Enlarged Cardiomediastinum, CM = Cardiomegaly, LO = Lung Opacity, LL = Lung Lesion, ED = Edema, CON = Consolidation, PNA = Pneumonia, ATL = Atelectasis, PTX = Pneumothorax, PE = Pleural Effusion, PO = Pleural Other, FX = Fracture, SD = Support Devices.
+Appendix G Additional Qualitative MXA examples on CXRs
+Figure G.1: Additional attention map visualizations after 25 training epochs are shown.
+Across CXRs with pneumonia and diverse acquisition geometries the MXA block consistently (i) reduces attention on bony anatomy and monitoring hardware, and (ii) reallocates it toward pleural regions containing abnormalities, corroborating the quantitative AUC gains reported in Tab.
+1.
+Bright yellow in the Naive/Improved panels = high attention; dark purple = low attention.
+The Delta panel shows the difference in attention: red indicates regions where MXA attends less than naive MHSA, blue where it attends more, and white means no change.
+Appendix H Multi-label EfficientViT design space
+Table H.1 lists the six principal variants of our multi-label EfficientViT architectures we defined.
+Each model family member differs in its stage-wise width ($C_{i}$), depth ($L_{i}$), and number of heads ($H_{i}$).
+Specifically, we break the network into three stages, and each stage’s dimensions are adjusted to control the overall capacity and computational cost.
+Resource limits restricted our experiments to the M5 configuration, which offers the best parameter–FLOP trade-off within that budget.
+Future work will explore the remaining variants.
+Table H.1: Multi-label EfficientViT architecture variants
+Model $\{C_{1},C_{2},C_{3}\}$$\{L_{1},L_{2},L_{3}\}$$\{H_{1},H_{2},H_{3}\}$
+EfficientViT-MultiLabel-M0 {64, 128, 192} {1, 2, 3} {4, 4, 4}
+EfficientViT-MultiLabel-M1 {128, 144, 192} {1, 2, 3} {2, 3, 3}
+EfficientViT-MultiLabel-M2 {128, 192, 224} {1, 2, 3} {4, 3, 2}
+EfficientViT-MultiLabel-M3 {128, 240, 320} {1, 2, 3} {4, 3, 4}
+EfficientViT-MultiLabel-M4 {128, 256, 384} {1, 2, 3} {4, 4, 4}
+EfficientViT-MultiLabel-M5 {192, 288, 384} {1, 3, 4} {3, 3, 4}
+Appendix I Additional quantitative plots
+Figure I.1: Mean AUC over three validation runs for baseline and proposed models; error bars denote 95% confidence intervals (±$\pm 2\sigma$).
+Across three runs, MXA + KD improves mean AUC drastically over the baseline, demonstrating robustness to random seed.
+Figure I.2: Box-plot comparison of per-epoch AUC for three runs (baseline vs.
+proposed).
+Every epoch shows higher and less variable AUC for MXA + KD, indicating both greater accuracy and consistency than the baseline.
+Appendix J Reported compute resources
+Table J.1: Compute Resources and Usage
+Resource Specification
+Lambda Labs Node
+   GPU 1$\times$ NVIDIA H100 (80 GB PCIe)
+   CPU 26 vCPUs
+   Memory 200 GiB RAM
+   Storage 1 TiB SSD
+   Number of experiments $\sim$120
+   Avg. compute time per experiment $\sim$10 hours
+   Total compute time $\sim$1,200 hours
+Google TPU
+   TPU 1$\times$ v4-8
+   Number of experiments $\sim$1
+   Avg. compute time per experiment $\sim$50 hours
+   Total compute time $\sim$ 50 hours